### v1

# An upper bound for the number of smooth values of a polynomial and its applications

Abstract.
We provide a new upper bound for the number of smooth values of a polynomial with integer coefficients. This improves Timofeev’s previous result unless the polynomial is a product of linear polynomials with integer coefficients. As an application, we fix an error in the proof of a result of Cassels which was used to prove that the Hurwitz zeta-function with algebraic irrational parameter has infinitely many zeros on the domain of convergence. We also apply the main result to a problem on primitive divisors of quadratic polynomials.
Key words and phrases:
smooth number, polynomial value, prime ideal, primitive divisor
2020 Mathematics Subject Classification:
Primary 11N32; Secondary 11N25
1. Introduction
1.1. Statement of the main results
Let $P^{+}(m)$ denote the largest prime factor of an integer $m$ with conventions $P^{+}(\pm 1)=1$ and $P^{+}(0)=+\infty$. An integer $m$ is said to be $y$-smooth or $y$-friable if $P^{+}(m)\leq y$ holds. The study of the distribution of smooth numbers is a classical topic in analytic number theory. Denote by $\Psi(x,y)$ the number of integers $n$ in the interval $[1,x]$ such that $n$ is $y$-smooth. In 1930, Dickman [5] showed that
$\Psi(x,x^{1/u})\sim\rho(u)x$
as $x\to\infty$ for every fixed $u>0$, where $\rho$ is the continuous solution of the system
$\begin{cases}\rho(u)=1&\text{for $0\leq u\leq 1$, }\\ u\rho^{\prime}(u)+\rho(u-1)=0&\text{for $u>1$.}\end{cases}$
The function $\rho$ is known today as the Dickman function. See the surveys [8, 10] for more results on the asymptotic behavior of $\Psi(x,y)$.
Let $f(t)\in\mathbb{Z}[t]$. Denote by $\Psi_{f}(x,y)$ the number of integers $n$ in $[1,x]$ such that $f(n)$ is $y$-smooth. Suppose that $f(t)$ has distinct irreducible factors over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. It was conjectured by Martin [15] that
(1.1) $\Psi_{f}(x,x^{1/u})\sim\rho(d_{1}u)\cdots\rho(d_{g}u)x$
as $x\to\infty$ for every fixed $u>0$. No result is known that (1.1) holds unconditionally for any non-linear polynomial $f(t)$ unless $u$ is too small to make it trivial. Then upper and lower bounds for $\Psi_{f}(x,y)$ have been studied. Hmyrova [11] proved that, if $f(t)$ is irreducible, then
$\Psi_{f}(x,x^{1/u})<c(f)\exp\left(-u\log{\frac{u}{e}}\right)x$
holds in the range $e\leq u\leq(\log{x})(\log\log{x})^{-1}$ for sufficiently large $x$, where $c(f)$ is a positive constant depending on the degree and coefficients of $f(t)$. For more general polynomials, Timofeev [18] obtained that
$\Psi_{f}(x,x^{1/u})<\frac{(g+\epsilon)^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}x$
holds in the range $1\leq u\leq\sqrt{(\log{x})(d+\epsilon)^{-1}}$ for sufficiently large $x$ and every $\epsilon>0$, where we put $d=d_{1}+\cdots+d_{g}$ under the same assumption as in (1.1). Here, and throughout this paper, $[u]$ denotes the largest integer not exceeding $u$. We omit the known results on lower bounds for $\Psi_{f}(x,y)$ because the theme of this paper is about the upper bounds. The statement of the main result is as follows.
Theorem 1.1.
Suppose that $f(t)\in\mathbb{Z}[t]$ has distinct irreducible factors over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. If $d=d_{1}+\cdots+d_{g}\geq 2$, then
(1.2) $\Psi_{f}(x,x^{1/u})<\gamma_{f}(u)\frac{g^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(x+O\left(\frac{x}{\log\log{x}}\right)\right)$
holds in the range $1\leq u\leq\sqrt{\log{x}}\,(\log\log{x})^{-1}$ for sufficiently large $x$, where $\gamma_{f}(u)$ is a positive real number which can be explicitly represented as
(1.3) $\gamma_{f}(u)=\frac{1}{2}+\frac{2g+1}{16du}+\sqrt{\frac{2g+1}{16du}+\left(\frac{2g+1}{16du}\right)^{2}}.$
The implied constant in (1.2) depends only on $f(t)$.
Suppose further that at least one of the irreducible factors of $f(t)$ is not a linear polynomial. Then we see that $g<d$ is satisfied, which implies
$\frac{2g+1}{16du}\leq\frac{2d-1}{16d}<\frac{1}{8}$
for any $u\geq 1$. In this case, we obtain
$\gamma_{f}(u)<\frac{1}{2}+\frac{1}{8}+\sqrt{\frac{1}{8}+\frac{1}{64}}=1.$
Hence, (1.2) improves on Timofeev’s upper bound unless $f(t)$ is a product of linear polynomials. Furthermore, if $f(t)$ is an irreducible polynomial of degree $d\geq 2$, then
(1.4) $\gamma_{f}(u)=\frac{1}{2}+\frac{3}{16du}+\sqrt{\frac{3}{16du}+\left(\frac{3}{16du}\right)^{2}}\leq\frac{19+\sqrt{105}}{32}=0.913...$
holds for any $u\geq 1$. This is also useful in applications of Theorem 1.1.
Theorem 1.1 is proved by induction on $m=[u]$. In Section 2, we overview the strategy of the proof and show some estimates in the case $m=1$. Then we explain the inductive process in Section 3. The proof of Theorem 1.1 is finally completed in Section 4.
1.2. Note to “Footnote to a note of Davenport and Heilbronn”
Let $\alpha$ be an arbitrary algebraic irrational number, and $K=\mathbb{Q}(\alpha)$ the algebraic field generated by $\alpha$ over the rationals. Let $\mathfrak{a}$ be the ideal denominator of the fractional ideal $(\alpha)$ of $K$. Then we see that $(n+\alpha)\mathfrak{a}$ is an integral ideal for every $n\in\mathbb{Z}$. Denote by $C_{\alpha}(x)$ the number of integers $n$ in $[1,x]$ such that there exists a prime ideal $\mathfrak{p}$ of $K$ with the conditions
(1.5) $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}\quad\text{and}\quad\mathfrak{p}\nmid\prod_{\begin{subarray}{c}1\leq m\leq x\\ m\neq n\end{subarray}}(m+\alpha)\mathfrak{a}.$
In Section 5, we prove the following result by applying Theorem 1.1.
Theorem 1.2.
Let $\alpha$ be an arbitrary algebraic irrational number. Then we have
$C_{\alpha}(x)\geq\left\{1-\frac{1}{2d}-\frac{3}{16d^{2}}-\frac{1}{d}\sqrt{\frac{3}{16d}+\left(\frac{3}{16d}\right)^{2}}\right\}x$
for sufficiently large $x$, where $d=\deg(\alpha)\geq 2$. In particular, the inequality
$C_{\alpha}(x)>0.543x$
holds for sufficiently large $x$.
Theorem 1.2 was essentially obtained by Cassels [2]. He applied this kind of result to study the zeros of the Hurwitz zeta-function $\zeta(s,\alpha)$ for $\operatorname{Re}(s)>1$, where $\alpha$ is an algebraic irrational number with $10<\alpha<$ 1. Indeed, he showed that $\zeta(s,\alpha)$ has infinitely many zeros in the strip $1<\operatorname{Re}(s)<1+\delta$ for any $\delta>0$, which extends the classical result of Davenport and Heilbronn [4]. However, the proof in [2] contains an error which cannot be easily fixed as explained in [16]. See Remark 5.5 for more details of the difference between the methods of [2] and this paper.
Remark that the formulation of Theorem 1.2 is due to Worley [20], and the original statement by Cassels in [2] is slightly different. We discuss how Cassels’ result is proved at the end of Section 5.
1.3. Primitive divisors of $n^{2}+b$
Another application of Theorem 1.1 is about the primitive divisors of quadratic polynomials. Let $(A_{n})$ be a sequence of integers. We say that an integer $d>1$ is a primitive divisor of $A_{n}$ if $d\mid A_{n}$ and $(d,A_{m})=1$ for all nonzero terms $A_{m}$ with $m<n$. Then we define 111 In the previous papers, $R_{b}(x)$ was denoted by $\rho_{b}(x)$, and usually $\rho_{1}(x)$ was abbreviated to $\rho(x)$. We change the symbols so as to keep that $\rho$ represents the Dickman function.
$R_{b}(x)=\#\left\{n\in\mathbb{Z}\cap[1,x]~{}\middle|~{}\textrm{$n^{2}+b$ has a primitive divisor}\right\}$
for $b\in\mathbb{Z}$. Suppose that $-b$ is not an integer square. Then Everest and Harman [6] conjectured that $R_{b}(x)\sim(\log{2})x$ as $x\to\infty$. They also proved
$0.5324x<R_{b}(x)<0.905x$
for sufficiently large $x$. Recently, Harman [9] sharpened the result for $R_{1}(x)$ to
$0.5377x<R_{1}(x)<0.86x.$
Li [14] also announced an even better upper bound $R_{1}(x)<0.8437x$. In this paper, we provide a new lower bound for $R_{b}(x)$.
Theorem 1.3.
Suppose that $-b$ is not an integer square. Then the inequality
$R_{b}(x)>0.543x$
holds for sufficiently large $x$.
The proof of Theorem 1.3 is presented in Section 6. It should be noted that the method for the proof is applicable only to the lower bound for $R_{b}(x)$, and we have no contribution to the upper bound.
For $b=1$, Theorem 1.3 is also related to a problem on arctangent relations which was originally studied by Todd [19]. For $n\in\mathbb{Z}_{\geq 1}$, we say that $\arctan n$ is reducible if there exist integers $f_{1},\ldots,f_{k}$ and $n_{1},\ldots,n_{k}$ with $1\leq n_{j}<n$ such that
$\arctan n=f_{1}\arctan n_{1}+\cdots+f_{k}\arctan n_{k}$
holds. For example, $\arctan 3$ is reducible by $\arctan 3=3\arctan 1-\arctan 2$. On the contrary, we say that $\arctan n$ is irreducible if it is not reducible. Then we define
$N(x)=\#\left\{n\in\mathbb{Z}\cap[1,x]~{}\middle|~{}\text{$\arctan n$ is irreducible}\right\}.$
Chowla and Todd [3] conjectured that $N(x)\sim(\log{2})x$ as $x\to\infty$. Kowalski [13] applied an equidistribution result on the roots of $X^{2}+1\equiv 0~{}(\bmod\,{p})$ to show that
$N(x)>(2-\epsilon+o(1))\frac{x}{\log{x}}$
as $x\to\infty$ for any $\epsilon>0$. Actually, it is not difficult to see that $N(x)=R_{1}(x)$ holds; see Lemmas 6.1 and 6.2. Hence we obtain the following result.
Theorem 1.4.
For sufficiently large $x$, we have $N(x)>0.543x$.
It would be worth noting that $N(x)<0.8437x$ also follows from Li’s work [14].
Acknowledgments
The author is especially grateful to Yuta Suzuki for sending a copy and translation of Timofeev’s paper [18]. Many discussions with him were very fruitful. The author would also like to thank Hiroshi Mikawa. Theorems 1.3 and 1.4 were motivated by his lecture at Hachiouji Number Theory Seminar 2024. The work of this paper was supported by JSPS Grant-in-Aid for Early-Career Scientists (Grant Number 24K16906).
2. Strategy and the initial estimates
We begin by a simple observation. Let $f(t)\in\mathbb{Z}[t]$ be any polynomial of positive degree, and $x,y,z\in\mathbb{R}$ with $x>z\geq 1$ and $y\geq 1$. Suppose that the inequality
(2.1) $\Psi_{f}(x,y)-\Psi_{f}(z,y)<V+\sqrt{\Psi_{f}(x,y)-\Psi_{f}(z,y)}\sqrt{W}$
holds with some parameters $V,W>0$. Then it can be easily seen that
(2.2) $\Psi_{f}(x,y)-\Psi_{f}(z,y)<V+\frac{W}{2}+\sqrt{VW+\frac{W^{2}}{4}}$
is satisfied. Therefore we would like to find suitable $V$ and $W$ such that (2.1) holds. Here, we notice that we may suppose that $f(t)\to\infty$ as $t\to\infty$ without loss of generality since $\Psi_{-f}(x,y)=\Psi_{f}(x,y)$ holds by definition. In this case, there exists a constant $T_{0}(f)\geq 2$ such that
(2.3) $f(t_{1})>f(t_{2})>1$
is valid for every $t_{1},t_{2}\in\mathbb{R}$ with $t_{1}>t_{2}>T_{0}(f)$. The following result is the initial step in the investigation of $V$ and $W$.
Proposition 2.1.
Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$. Then inequality (2.1) holds with
$V$ $=\frac{1}{\log{f(z)}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\Lambda(k)\sum_{\begin{subarray}{c}z<n\leq x\\ k\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1,$
$W$ $=\frac{1}{(\log{f(z)})^{2}}\sum_{\begin{subarray}{c}1\leq k_{1},k_{2}\leq f(x)\\ P^{+}(k_{1})\leq\sqrt{y}\\ P^{+}(k_{2})\leq\sqrt{y}\end{subarray}}\Lambda(k_{1})\Lambda(k_{2})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1,$
where $[k_{1},k_{2}]$ denotes the largest common multiple of $k_{1}$ and $k_{2}$.
Proof.
The idea of the proof is borrowed from Cassels [2]. It is deduced from (2.3) that $\log{f(n)}>\log{f(z)}>0$ for $n>z>T_{0}(f)$. Thus we have
$\Psi_{f}(x,y)-\Psi_{f}(z,y)=\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}1<\frac{1}{\log{f(z)}}\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\log{f(n)}$
by the definition of $\Psi_{f}(x,y)$. Considering the prime factorization of $f(n)$, we obtain
$\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\log{f(n)}=\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}p,v\\ p^{v}\mid f(n)\end{subarray}}\log{p},$
where $\sum_{p,v}$ stands for the double sum over prime numbers $p$ and integers $v\geq 1$. Let $\Lambda(n)$ denote as usual the von Mangoldt function. Then we have
$\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}p,v\\ p^{v}\mid f(n)\end{subarray}}\log{p}$ $=\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ P^{+}(k)\leq y\end{subarray}}\Lambda(k)$
$=\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\Lambda(k)+\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ P^{+}(k)\leq\sqrt{y}\end{subarray}}\Lambda(k).$
From the above, we see that the inequality
(2.4) $\Psi_{f}(x,y)-\Psi_{f}(z,y)$ $<\frac{1}{\log{f(z)}}\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\Lambda(k)$
$\quad+\frac{1}{\log{f(z)}}\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ P^{+}(k)\leq\sqrt{y}\end{subarray}}\Lambda(k)$
holds. The first sum of the right-hand side of (2.4) is evaluated as
(2.5) $\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\Lambda(k)=\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\Lambda(k)\sum_{\begin{subarray}{c}z<n\leq x\\ k\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
by changing the order of the summations. As for the second sum of the right-hand side of (2.4), we apply the Cauchy–Schwarz inequality to deduce
$\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ P^{+}(k)\leq\sqrt{y}\end{subarray}}\Lambda(k)\leq\sqrt{\Psi_{f}(x,y)-\Psi_{f}(z,y)}\sqrt{\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\Bigg{(}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ P^{+}(k)\leq\sqrt{y}\end{subarray}}\Lambda(k)\Bigg{)}^{2}}.$
Furthermore, we have the identity
$\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\Bigg{(}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ P^{+}(k)\leq\sqrt{y}\end{subarray}}\Lambda(k)\Bigg{)}^{2}=\sum_{\begin{subarray}{c}1\leq k_{1},k_{2}\leq f(x)\\ P^{+}(k_{1})\leq\sqrt{y}\\ P^{+}(k_{2})\leq\sqrt{y}\end{subarray}}\Lambda(k_{1})\Lambda(k_{2})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
by expanding the square and changing the order of summations. Hence we arrive at
(2.6) $\sum_{\begin{subarray}{c}z<n\leq x\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq k\leq f(x)\\ k\mid f(n)\\ P^{+}(k)\leq\sqrt{y}\end{subarray}}\Lambda(k)$
$\leq\sqrt{\Psi_{f}(x,y)-\Psi_{f}(z,y)}\sqrt{\sum_{\begin{subarray}{c}1\leq k_{1},k_{2}\leq f(x)\\ P^{+}(k_{1})\leq\sqrt{y}\\ P^{+}(k_{2})\leq\sqrt{y}\end{subarray}}\Lambda(k_{1})\Lambda(k_{2})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1}.$
Then the desired result follows by inserting (2.5) and (2.6) to inequality (2.4). ∎
3. An inductive process
By Proposition 2.1, the next task to deduce an upper bound of $\Psi_{f}(x,y)-\Psi_{f}(z,y)$ is evaluating the sum
$S(\kappa)=\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
with $\kappa=k$ and $\kappa=[k_{1},k_{2}]$. For every $\kappa\in\mathbb{Z}_{\geq 1}$, we define $\omega_{f}(\kappa)$ as the number of the zeros of the polynomial $f(t)$ in the residue class ring $\mathbb{Z}/\kappa\mathbb{Z}$, that is,
(3.1) $\omega_{f}(\kappa)=\#\left\{u~{}(\bmod\,{\kappa})~{}\middle|~{}f(u)\equiv 0~{}(\bmod\,{\kappa})\right\}.$
Recall that $f(j\kappa+u)\equiv f(u)~{}(\bmod\,{\kappa})$ for any $j\in\mathbb{Z}$. Therefore we obtain
(3.2) $S(\kappa)\leq\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\end{subarray}}1=\sum_{\begin{subarray}{c}u\pmod{\kappa}\\ f(u)\equiv 0\pmod{\kappa}\end{subarray}}\sum_{\begin{subarray}{c}z<n\leq x\\ n\equiv u\pmod{\kappa}\end{subarray}}1\leq\omega_{f}(\kappa)\left(\frac{x-z}{\kappa}+1\right),$
but this estimate is not good when $y$ is small since we have ignored the condition $P^{+}(f(n))\leq y$ in the first inequality. The purpose of this section is to evaluate $S(\kappa)$ by using $S(\kappa\lambda)$ for $\lambda\geq 1$, which enables us to refine inductively the estimates of $V$ and $W$ such that (2.1) holds.
Lemma 3.1.
Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$. Suppose that $f(z)>x$ is satisfied, and put $h=x-z$. Then we have
(3.3) $\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$ $<\frac{1}{\log{\frac{f(z)}{x}}}\sum_{\begin{subarray}{c}1\leq\lambda\leq h/\kappa\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
$\quad+\frac{1}{\log{\frac{f(z)}{x}}}\sum_{\begin{subarray}{c}h/\kappa<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(\kappa\lambda)$
for any $\kappa\in\mathbb{Z}$ with $1\leq\kappa\leq h$.
Proof.
Since we have $\kappa\leq h<x$ and $f(z)>x$, we find that $\log{\frac{f(n)}{\kappa}}>\log{\frac{f(z)}{x}}>0$ for $n>z>T_{0}(f)$ by (2.3). Therefore we obtain the inequality
$\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1<\frac{1}{\log{\frac{f(z)}{x}}}\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}\log{\frac{f(n)}{\kappa}}.$
Considering the prime factorization of $f(n)/\kappa$, we have
$\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}\log{\frac{f(n)}{\kappa}}=\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq\lambda\leq f(x)\\ \lambda\mid f(n)/\kappa\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)$
by the same argument as in the proof of Proposition 2.1. Furthermore, we have
$\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}\sum_{\begin{subarray}{c}1\leq\lambda\leq f(x)\\ \lambda\mid f(n)/\kappa\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)$ $=\sum_{\begin{subarray}{c}1\leq\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
$=\sum_{\begin{subarray}{c}1\leq\lambda\leq h/\kappa\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1+\sum_{\begin{subarray}{c}h/\kappa<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
by changing the order of summations and dividing the sum over $\lambda$. The sum over $\lambda$ with $1\leq\lambda\leq h/\kappa$ yields the first term of the right-hand side of (3.3). On the other hand, the sum over remaining $\lambda$ is estimated as
$\sum_{\begin{subarray}{c}h/\kappa<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1\leq\sum_{\begin{subarray}{c}h/\kappa<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(\kappa\lambda)$
since the number of integers $n$ with $z<n\leq x$ and $\kappa\lambda\mid f(n)$ is less than $\omega_{f}(\kappa\lambda)$ due to $h=x-z<\kappa\lambda$. Hence we obtain the second term of the right-hand side of (3.3), and the proof is completed. ∎
Lemma 3.1 was developed by Timofeev [18]. He used this result implicitly in his study on the upper bound for $\Psi_{f}(x,y)$. In this paper, we apply Lemma 3.1 to refine Proposition 2.1 as follows.
Proposition 3.2.
Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$. Suppose that $f(z)>x$ is satisfied, and put $h=x-z$. Then inequality (2.1) holds with
(3.4) $V=V_{m}^{+}+\sum_{i=1}^{m}V_{i}^{-}\quad\text{and}\quad W=W_{m}^{+}+\sum_{i=1}^{m}W_{i}^{-}$
for every $m\in\mathbb{Z}_{\geq 1}$, where $V_{m}^{+}$, $W_{m}^{+}$, $V_{i}^{-}$, $W_{i}^{-}$ for $i=1,\ldots,m$ are defined as follows. Firstly, we define $V_{m}^{+}$ and $W_{m}^{+}$ as
$V_{m}^{+}$ $=\frac{1}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}z<n\leq x\\ k_{1}\cdots k_{m}\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1,$
$W_{m}^{+}$ $=\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}$
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
with the convention $k_{3}\cdots k_{m+1}=1$ for $m=1$. Secondly, we define $V_{i}^{-}$ as
$V_{i}^{-}=\frac{1}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{i-1}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i}\geq 1\\ k_{1}\cdots k_{i-1}\leq h,~{}\frac{h}{k_{1}\cdots k_{i-1}}<k_{i}\leq f(x)\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i})\omega_{f}(k_{1}\cdots k_{i}),$
for every $1\leq i\leq m$ with the convention $k_{1}\cdots k_{i-1}=1$ for $i=1$. Lastly, we define
$W_{i}^{-}$ $=\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{i-1}}$
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{i}\leq h,~{}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{i}}<k_{i+1}\leq f(x)\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{i+1})$
for every $2\leq i\leq m$ with the convention $k_{3}\cdots k_{i}=1$ for $i=2$, and
$W_{1}^{-}=\frac{1}{(\log{f(z)})^{2}}\sum_{\begin{subarray}{c}1\leq k_{1},k_{2}\leq f(x)\\ [k_{1},k_{2}]>h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\Lambda(k_{2})\omega_{f}([k_{1},k_{2}]).$
Proof.
The proof is by induction on $m$. The result for $m=1$ can be directly deduced from Proposition 2.1. Hence it suffices to show that the inequalities
(3.5) $V_{m}^{+}<V_{m+1}^{+}+V_{m+1}^{-}\quad\text{and}\quad W_{m}^{+}<W_{m+1}^{+}+W_{m+1}^{-}$
hold for every $m\in\mathbb{Z}_{\geq 1}$ since they imply
$V_{m}^{+}+\sum_{i=1}^{m}V_{i}^{-}<V_{m+1}^{+}+\sum_{i=1}^{m+1}V_{i}^{-}\quad\text{and}\quad W_{m}^{+}+\sum_{i=1}^{m}W_{i}^{-}<W_{m+1}^{+}+\sum_{i=1}^{m+1}W_{i}^{-}.$
We apply Lemma 3.1 with $\kappa=k_{1}\cdots k_{m}$ to derive
$V_{m}^{+}$ $<\frac{1}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{k_{1}\cdots k_{m}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ k_{1}\cdots k_{m}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
$\quad+\frac{1}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}\frac{h}{k_{1}\cdots k_{m}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
Replacing $\lambda$ by $k_{m+1}$, we see that the first and second terms of the right-hand side are equal to $V_{m+1}^{+}$ and $V_{m+1}^{-}$, respectively. Hence we obtain inequality (3.5) for $V_{m}^{+}$. Similarly, we apply Lemma 3.1 with $\kappa=[k_{1},k_{2}]k_{3}\cdots k_{m+1}$ to derive
$W_{m}^{+}<\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
$\quad+\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
In this case, we replace $\lambda$ with $k_{m+2}$. Then we see that the first and second terms of the right-hand side are equal to $W_{m+1}^{+}$ and $W_{m+1}^{-}$, respectively. Then inequality (3.5) follows also for $W_{m}^{+}$. ∎
4. Proof of Theorem 1.1
Before proceeding to the completion of the proof of Theorem 1.1, we notice that we may suppose that the polynomial $f(t)$ is a product of distinct irreducible factors $f_{1}(t),\ldots,f_{g}(t)$ over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. Indeed, if $f(t)$ is factorized in general as
$f(t)=Af_{1}(t)^{e_{1}}\cdots f_{g}(t)^{e_{g}}$
with $A,e_{1},\ldots,e_{g}\in\mathbb{Z}$ such that $A\neq 0$ and $e_{j}\geq 1$, then it is deduced from the definition of $\Psi_{f}(x,y)$ that $\Psi_{\operatorname{rad}(f)}(x,y)=\Psi_{f}(x,y)$ for $y>|A|$, where we define
$\operatorname{rad}(f(t))=f_{1}(t)\cdots f_{g}(t).$
Then we prove the following lemma on sums involving $\omega_{f}(k)$ defined as (3.1).
Lemma 4.1.
Suppose that $f(t)\in\mathbb{Z}[t]$ is a product of distinct irreducible factors $f_{1}(t),\ldots,f_{g}(t)$ over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. Then we have
(4.1) $\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)$ $\leq g\log{y}+O(1),$
(4.2) $\sum_{\begin{subarray}{c}1\leq k\leq x\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)$ $\leq\frac{g}{2}\log{y}+O(1),$
(4.3) $\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{(2\log{k}-\Lambda(k))\Lambda(k)}{k}\omega_{f}(k)$ $\leq\frac{g}{2}(\log{y})^{2}+O(\log{y}),$
(4.4) $\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\Lambda(k)\omega_{f}(k)$ $=O\left(y\frac{\log{x}}{\log{y}}\right)$
for any $x,y\geq 2$, where all implied constants depend only on $f(t)$.
Proof.
Let $\Delta_{f}$ denote the absolute value of the discriminant of $f(t)$. Note that $\Delta_{f}$ is a positive integer by the assumption on $f(t)$. Then we write the prime factorization of $\Delta_{f}$ as
$\Delta_{f}=\prod_{p}p^{\theta(p)}$
with $\theta(p)\in\mathbb{Z}$ such that $\theta(p)=0$ for all but finitely many of the prime numbers $p$. Huxley [12] proved the bound $\omega_{f}(p^{v})\leq dp^{\theta(p)/2}$ with $d=d_{1}+\cdots+d_{g}$. Thus we obtain the uniform bound
(4.5) $\omega_{f}(p^{v})\leq d\sqrt{\Delta_{f}}.$
First, we consider estimates (4.1) and (4.2). Since the von Mangoldt function $\Lambda(n)$ is supported on prime powers, we have
(4.6) $\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)$ $=\sum_{p\leq y}\sum_{\begin{subarray}{c}v\geq 1\\ p^{v}\leq x\end{subarray}}\frac{\log{p}}{p^{v}}\omega_{f}(p^{v})$
$=\sum_{p\leq y}\frac{\log{p}}{p}\omega_{f}(p)+\sum_{p\leq y}\sum_{\begin{subarray}{c}v\geq 2\\ p^{v}\leq x\end{subarray}}\frac{\log{p}}{p^{v}}\omega_{f}(p^{v}).$
Using (4.5), we see that the latter term is estimated as
(4.7) $\sum_{p\leq y}\sum_{\begin{subarray}{c}v\geq 2\\ p^{v}\leq x\end{subarray}}\frac{\log{p}}{p^{v}}\omega_{f}(p^{v})\leq d\sqrt{\Delta_{f}}\sum_{p}\sum_{v\geq 2}\frac{\log{p}}{p^{v}}\ll 1$
with the implied constant depending only on $d$ and $\Delta_{f}$. Note that we have
$\omega_{f}(p)\leq\omega_{f_{1}}(p)+\cdots+\omega_{f_{g}}(p)$
since $p\mid f(u)$ implies $p\mid f_{j}(u)$ for some $j\in\{1,\ldots,g\}$. Furthermore, it is classically known (see [17] for example) that
(4.8) $\sum_{p\leq y}\frac{\log{p}}{p}\omega_{f_{j}}(p)=\log{y}+O(1)$
for every $j\in\{1,\ldots,g\}$ with the implied constant depends only on $K_{j}=\mathbb{Q}(\alpha_{j})$, where $\alpha_{j}\in\mathbb{C}$ is a solution to the equation $f_{j}(t)=0$. Hence we have
(4.9) $\sum_{p\leq y}\frac{\log{p}}{p}\omega_{f}(p)\leq g\log{y}+O(1),$
where the implied constant depends at most on $f(t)$. Inserting (4.7) and (4.9) to (4.6), we obtain (4.1). One can obtain (4.2) along the same line, and we omit the proof. Next, we prove estimate (4.3) as follows. We have
$\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{(2\log{k}-\Lambda(k))\Lambda(k)}{k}\omega_{f}(k)$ $=\sum_{p\leq y}\sum_{\begin{subarray}{c}v\geq 1\\ p^{v}\leq x\end{subarray}}\frac{(2v-1)(\log{p})^{2}}{p^{v}}\omega_{f}(p^{v})$
$=\sum_{p\leq y}\frac{(\log{p})^{2}}{p}\omega_{f}(p)+O(1)$
in a way similar to (4.6) and (4.7). Here, one can check that
$\sum_{p\leq y}\frac{(\log{p})^{2}}{p}\omega_{f_{j}}(p)=\frac{1}{2}(\log{y})^{2}+O(\log{y})$
holds for every $j\in\{1,\ldots,g\}$ by the partial summation using formula (4.8). Hence (4.3) follows by the same argument as before. Finally, we obtain
$\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\Lambda(k)\omega_{f}(k)=\sum_{p\leq y}\sum_{\begin{subarray}{c}v\geq 1\\ p^{v}\leq x\end{subarray}}(\log{p})\omega_{f}(p^{v})$ $\leq d\sqrt{\Delta_{f}}\sum_{p\leq y}\left[\frac{\log{x}}{\log{p}}\right](\log{p})$
$\ll(\log{x})\sum_{p\leq y}1,$
where the implied constant depends only on $d$ and $\Delta_{f}$. Then we deduce that estimate (4.4) holds by the prime number theorem. ∎
Note that the function $\omega_{f}(k)$ is multiplicative but is not completely multiplicative in general. Hence we do not expect that $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds unless $k_{1},\ldots,k_{m}$ are pairwise coprime. 222 It seems to be misunderstood often in [18] that the identity $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds for any $k_{1},\ldots,k_{m}\in\mathbb{Z}_{\geq 1}$. One can fix the errors with the same argument presented in this paper using Lemma 4.2. Then we use the following lemma which helps us estimate multiple sums involving $\omega_{f}(k_{1}\cdots k_{m})$.
Lemma 4.2.
Suppose that $f(t)\in\mathbb{Z}[t]$ is a product of distinct irreducible factors $f_{1}(t),\ldots,f_{g}(t)$ over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. Put $d=d_{1}+\cdots+d_{g}$, and denote by $\Delta_{f}$ the absolute value of the discriminant of $f(t)$. Then we have
$\omega_{f}(p_{1}^{v_{1}}\cdots p_{m}^{v_{m}})\leq(d\sqrt{\Delta_{f}})^{m-|S|}\prod_{j\in S}\omega_{f}(p_{j}^{v_{j}})$
for any prime numbers $p_{1},\ldots,p_{m}$ and any integers $v_{1},\ldots,v_{m}\geq 1$, where $S$ is the subset of $\{1,\ldots,m\}$ such that $j\in S$ if and only if $p_{j}\nmid\Delta_{f}$.
Proof.
Put $p_{1}^{v_{1}}\cdots p_{m}^{v_{m}}=AB$ with $A=\prod_{j\in S}p_{j}^{v_{j}}$ and $B=\prod_{j\notin S}p_{j}^{v_{j}}$. Since the function $\omega_{f}(k)$ is multiplicative, we have
$\omega_{f}(p_{1}^{v_{1}}\cdots p_{m}^{v_{m}})=\omega_{f}(A)\omega_{f}(B)$
by noting that $(A,B)=1$. Here, and throughout this paper, $(A,B)$ denotes the greatest common divisor of $A$ and $B$. Remark that the result is trivial if $\omega_{f}(A)=0$. Thus we suppose that $\omega_{f}(A)\neq 0$ below. Let $S^{\prime}$ be a minimal subset of $S$ such that $\{p_{j}\mid j\in S\}=\{p_{j}\mid j\in S^{\prime}\}$, and rewrite the integer $A$ as $A=\prod_{j\in S^{\prime}}p_{j}^{w_{j}}$. In this case, the primes $p_{j}$ for $j\in S^{\prime}$ are distinct. Therefore we have
$\omega_{f}(A)=\prod_{j\in S^{\prime}}\omega_{f}(p_{j}^{w_{j}})$
by the multiplicativity of $\omega_{f}(k)$. Furthermore, we see that $\omega_{f}(p_{j}^{w_{j}})=\omega_{f}(p_{j})$ holds for any $j\in S^{\prime}$ by Hensel’s lemma. Thus $\omega_{f}(A)$ is evaluated as
$\omega_{f}(A)=\prod_{j\in S^{\prime}}\omega_{f}(p_{j})\leq\prod_{j\in S}\omega_{f}(p_{j})$
since $\omega_{f}(p_{j})\geq 1$ is valid for any $j\in S$ by the assumption that $\omega_{f}(A)\neq 0$. Applying Hensel’s lemma again, we have $\omega_{f}(p_{j})=\omega_{f}(p_{j}^{v_{j}})$ for any $j\in S$. Hence we conclude that the inequality
(4.10) $\omega_{f}(A)\leq\prod_{j\in S}\omega_{f}(p_{j}^{v_{j}})$
is satisfied. Let $S^{\prime\prime}$ be a maximal subset of $\{1,\ldots,m\}$ with $S\subset S^{\prime\prime}$ such that $\{p_{j}\mid j\notin S\}=\{p_{j}\mid j\notin S^{\prime\prime}\}$, and rewrite the integer $B$ as $B=\prod_{j\notin S^{\prime\prime}}p_{j}^{w_{j}}$. In this case, we derive
(4.11) $\omega_{f}(B)=\prod_{j\notin S^{\prime\prime}}\omega_{f}(p_{j}^{w_{j}})\leq(d\sqrt{\Delta_{f}})^{m-|S^{\prime\prime}|}\leq(d\sqrt{\Delta_{f}})^{m-|S|}$
by using (4.5). Combining (4.10) and (4.11), we obtain the desired result. ∎
We are now ready to evaluate $V_{m}^{+}$, $W_{m}^{+}$, and $V_{i}^{-}$, $W_{i}^{-}$ for $i=1,\ldots,m$ defined in Proposition 3.2. Firstly, we consider the upper bounds for $V_{m}^{+}$ and $V_{i}^{-}$.
Proposition 4.3.
Suppose that $f(t)\in\mathbb{Z}[t]$ is a product of distinct irreducible factors $f_{1}(t),\ldots,f_{g}(t)$ over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. Let $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 2$. Suppose further that $f(t)\to\infty$ as $t\to\infty$, and that $f(z)>x$. Put $h=x-z$. Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
$V_{m}^{+}$ $\leq\frac{hg^{m}(\log{y}+O(1))^{m}}{2\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}+\frac{A(f)^{m}(\log{x})^{m}\left(\frac{y}{\log{y}}\right)^{m}}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m-1}},$
$V_{i}^{-}$ $\leq\frac{A(f)^{i}(\log{x})^{i-1}\log{f(x)}\left(\frac{y}{\log{y}}\right)^{i}}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{i-1}},$
where $V_{m}^{+}$ and $V_{i}^{-}$ are as in Proposition 3.2, and the implied constant depends only on $f(t)$.
Proof.
Using (3.2) with $\kappa=k_{1}\cdots k_{m}$, we have the inequality
$\sum_{\begin{subarray}{c}z<n\leq x\\ k_{1}\cdots k_{m}\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1\leq\omega_{f}(k_{1}\cdots k_{m})\left(\frac{h}{k_{1}\cdots k_{m}}+1\right).$
Then $V_{m}^{+}$ is evaluated as
(4.12) $V_{m}^{+}$ $\leq\frac{h}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m})}{k_{1}\cdots k_{m}}\omega_{f}(k_{1}\cdots k_{m})$
$\quad+\frac{1}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\omega_{f}(k_{1}\cdots k_{m}).$
Let $S\subset\{1,\ldots,m\}$. Then we divide the first sum of the right-hand side of (4.12) into those over $k_{1},\ldots,k_{m}\geq 1$ such that $(k_{j},\Delta_{f})=1$ if $j\in S$ and $(k_{j},\Delta_{f})>1$ otherwise, where $\Delta_{f}$ is the absolute value of the discriminant of $f(t)$. We have
(4.13) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\end{subarray}}=\sum_{S\subset\{1,\ldots,m\}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}$
$=\sum_{n=1}^{m}\sum_{\begin{subarray}{c}S\subset\{1,\ldots,m\}\\ |S|=n\\ 1\in S\end{subarray}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}+\sum_{n=0}^{m-1}\sum_{\begin{subarray}{c}S\subset\{1,\ldots,m\}\\ |S|=n\\ 1\notin S\end{subarray}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}$
by dividing further the cases into $1\in S$ or not. Suppose that $|S|=n$ and $1\in S$ hold. Since the von Mangoldt function $\Lambda(n)$ is supported on prime powers, we deduce from Lemma 4.2 that
$\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m})}{k_{1}\cdots k_{m}}\omega_{f}(k_{1}\cdots k_{m})$
$\leq\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m})}{k_{1}\cdots k_{m}}(d\sqrt{\Delta_{f}})^{m-n}\prod_{j\in S}\omega_{f}(k_{j})$
$\leq\sum_{\begin{subarray}{c}1\leq k\leq x\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\left(\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\right)^{n-1}\left(d\sqrt{\Delta_{f}}\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\\ (k,\Delta_{f})>1\end{subarray}}\frac{\Lambda(k)}{k}\right)^{m-n}$
by recalling that $h=x-z<x$. Applying (4.1) and (4.2) of Lemma 4.1, we have
$\sum_{\begin{subarray}{c}1\leq k\leq x\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\left(\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\right)^{n-1}$
$\leq\left(\frac{g}{2}\log{y}+O(1)\right)\left(g\log{y}+O(1)\right)^{n-1}=\frac{1}{2}\left(g\log{y}+O(1)\right)^{n},$
where the implied constants depend only on $f(t)$. Then we consider the sum over $k$ with $(k,\Delta_{f})>1$. We may suppose that $k$ varies over prime powers since $\Lambda(k)=0$ otherwise. Furthermore, if $k=p^{v}$ satisfies $(k,\Delta_{f})>1$, then $p\mid\Delta_{f}$ must be satisfied. Therefore we obtain
$\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\\ (k,\Delta_{f})>1\end{subarray}}\frac{\Lambda(k)}{k}=\sum_{\begin{subarray}{c}p\leq y\\ p\mid\Delta_{f}\end{subarray}}\sum_{\begin{subarray}{c}v\geq 1\\ p^{v}\leq x\end{subarray}}\frac{\log{p}}{p^{v}}\leq\sum_{p\mid\Delta_{f}}\frac{\log{p}}{p-1}\leq\log{\Delta_{f}}.$
Hence we arrive at the estimate
(4.14) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m})}{k_{1}\cdots k_{m}}\omega_{f}(k_{1}\cdots k_{m})$
$\leq\frac{1}{2}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
if $|S|=n$ and $1\in S$. On the other hand, if we suppose that $|S|=n$ and $1\notin S$, then we have
(4.15) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m})}{k_{1}\cdots k_{m}}\omega_{f}(k_{1}\cdots k_{m})$
$\leq\left(\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\right)^{n}\left(d\sqrt{\Delta_{f}}\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\\ (k,\Delta_{f})>1\end{subarray}}\frac{\Lambda(k)}{k}\right)^{m-n}$
$\leq\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
in a similar way. Here, the number of the subsets $S\subset\{1,\ldots,m\}$ with $|S|=n$ and $1\in S$ equals to $\binom{m-1}{n-1}$, and the number of $S\subset\{1,\ldots,m\}$ with $|S|=n$ and $1\notin S$ equals to $\binom{m-1}{n}$. Therefore, by (4.13), (4.14), and (4.15), we derive
(4.16) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m})}{k_{1}\cdots k_{m}}\omega_{f}(k_{1}\cdots k_{m})$
$\leq\sum_{n=1}^{m}\binom{m-1}{n-1}\frac{1}{2}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
$\quad+\sum_{n=0}^{m-1}\binom{m-1}{n}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
$=\frac{1}{2}\sum_{n=0}^{m}\binom{m}{n}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
$\quad+\frac{1}{2}\sum_{n=0}^{m-1}\binom{m-1}{n}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
$=\frac{1}{2}(g\log{y}+O(1))^{m}+\frac{1}{2}d\sqrt{\Delta_{f}}\log{\Delta_{f}}(g\log{y}+O(1))^{m-1}$
$=\frac{1}{2}g^{m}(\log{y}+O(1))^{m},$
where the implied constants depend only on $f(t)$. Next, we consider the second sum of the right-hand side of (4.12). We have
$\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\\ (k,\Delta_{f})>1\end{subarray}}\Lambda(k)=\sum_{\begin{subarray}{c}p\leq y\\ p\mid\Delta_{f}\end{subarray}}\sum_{\begin{subarray}{c}v\geq 1\\ p^{v}\leq x\end{subarray}}\log{p}\leq\sum_{p\mid\Delta_{f}}\left[\frac{\log{x}}{\log{p}}\right]\log{p}\ll\log{x},$
where the implied constant depends only on $\Delta_{f}$. Using this and (4.4) of Lemma 4.1 in place of (4.1) and (4.2), we also obtain
(4.17) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\omega_{f}(k_{1}\cdots k_{m})$
$=\sum_{n=0}^{m}\sum_{\begin{subarray}{c}S\subset\{1,\ldots,m\}\\ |S|=n\end{subarray}}\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\\ (k_{j},\Delta_{f})=1\Leftrightarrow j\in S\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\omega_{f}(k_{1}\cdots k_{m})$
$\leq\sum_{n=0}^{m}\binom{m}{n}\left(C_{1}(f)y\frac{\log{x}}{\log{y}}\right)^{n}\left(d\sqrt{\Delta_{f}}C_{2}(f)\log{x}\right)^{m-n}$
$=\left(C_{1}(f)y\frac{\log{x}}{\log{y}}+d\sqrt{\Delta_{f}}C_{2}(f)\log{x}\right)^{m}$
$\leq A(f)^{m}(\log{x})^{m}\left(\frac{y}{\log{y}}\right)^{m},$
where $C_{i}(f)$ and $A(f)$ are positive constants depending at most on $f(t)$. As a result, the desired estimate for $V_{m}^{+}$ follows from (4.12). As for $V_{i}^{-}$, we have
$\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i}\geq 1\\ k_{1}\cdots k_{i-1}\leq h,~{}\frac{h}{k_{1}\cdots k_{i-1}}<k_{i}\leq f(x)\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i})\omega_{f}(k_{1}\cdots k_{i})$
$\leq A(f)^{i}(\log{x})^{i-1}\log{f(x)}\left(\frac{y}{\log{y}}\right)^{i}$
for every $i=1,\ldots,m$ along the same line as (4.17). This yields the desired estimate for $V_{i}^{-}$. From the above, the proof of the results are completed. ∎
Secondly, we consider the upper bounds for $W_{m}^{+}$ and $W_{i}^{-}$.
Proposition 4.4.
Suppose that $f(t)\in\mathbb{Z}[t]$ satisfies the same conditions as in Proposition 4.3. Let $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 2$. Put $h=x-z$. Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
$W_{m}^{+}$ $\leq\frac{hg^{m}(2g+1)(\log{y}+O(1))^{m+1}}{8(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}+\frac{A(f)^{m+1}(\log{x})^{m+1}\left(\frac{y}{\log{y}}\right)^{m+1}}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}},$
$W_{i}^{-}$ $\leq\frac{A(f)^{i+1}(\log{x})^{i-1}(\log{f(x)})^{2}\left(\frac{y}{\log{y}}\right)^{i+1}}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{i-1}}$
where $W_{m}^{+}$ and $W_{i}^{-}$ are as in Proposition 3.2, and the implied constant depends only on $f(t)$.
Proof.
Using (3.2) with $\kappa=[k_{1},k_{2}]k_{3}\cdots k_{m}$, we obtain
$W_{m}^{+}$ $\leq\frac{h}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}$
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
$\quad+\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}$
×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
by the definition of $W_{m}^{+}$. We have
$\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}=\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\\ (k_{1},k_{2})=1\end{subarray}}+\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\\ (k_{1},k_{2})>1\end{subarray}}.$
Then we divide the sums into those over $k_{1},\ldots,k_{m+1}\geq 1$ such that $(k_{j},\Delta_{f})=1$ if $j\in S$ and $(k_{j},\Delta_{f})>1$ otherwise, where $S$ is a subset of $\{1,\ldots,m+1\}$. Recall that $[k_{1},k_{2}]k_{3}\cdots k_{m+1}=k_{1}k_{2}k_{3}\cdots k_{m+1}$ holds if $(k_{1},k_{2})=1$. Therefore we obtain
(4.18) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\\ (k_{1},k_{2})=1\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
$\leq\sum_{n=2}^{m}\binom{m-2}{n-2}\frac{1}{4}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
$\quad+2\sum_{n=1}^{m-1}\binom{m-2}{n-1}\frac{1}{2}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
$\quad+\sum_{n=0}^{m-2}\binom{m-2}{n}\left(g\log{y}+O(1)\right)^{n}\left(d\sqrt{\Delta_{f}}\log{\Delta_{f}}\right)^{m-n}$
$=\frac{1}{4}g^{m+1}(\log{y}+O(1))^{m+1}$
by an argument similar to (4.16), where we divide the cases into the followings:
  • $(\mathrm{a})$
    $1,2\in S$;
  • $(\mathrm{b})$
    $1\in S$ and $2\notin S$, or $1\notin S$ and $2\in S$;
  • $(\mathrm{c})$
    $1,2\notin S$.
On the other hand, the sum over $k_{1},\ldots,k_{m+1}$ such that $(k_{1},k_{2})>1$ is estimated as follows. We have
$\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\\ (k_{1},k_{2})>1\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
$=\sum_{\begin{subarray}{c}\kappa,k_{3},\ldots,k_{m+1}\geq 1\\ \kappa k_{3}\cdots k_{m+1}\leq h\\ P^{+}(\kappa)\leq\sqrt{y}\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\sum_{\begin{subarray}{c}k_{1},k_{2}\geq 1\\ [k_{1},k_{2}]=\kappa\\ (k_{1},k_{2})>1\end{subarray}}\frac{\Lambda(k_{1})\Lambda(k_{2})}{[k_{1},k_{2}]}\frac{\Lambda(k_{3})\cdots\Lambda(k_{m+1})}{k_{3}\cdots k_{m+1}}\omega_{f}(\kappa k_{3}\cdots k_{m+1}).$
Furthermore, since the von Mangoldt function $\Lambda(n)$ is supported on prime powers, the inner sum is calculated as
$\sum_{\begin{subarray}{c}k_{1},k_{2}\geq 1\\ [k_{1},k_{2}]=\kappa\\ (k_{1},k_{2})>1\end{subarray}}\frac{\Lambda(k_{1})\Lambda(k_{2})}{[k_{1},k_{2}]}$ $=\frac{\Lambda(\kappa)}{\kappa}\sum_{k_{1}\mid\kappa}\Lambda(k_{1})+\frac{\Lambda(\kappa)}{\kappa}\sum_{k_{2}\mid\kappa}\Lambda(k_{2})-\frac{\Lambda(\kappa)^{2}}{\kappa}$
$=\frac{(2\log{\kappa}-\Lambda(\kappa))\Lambda(\kappa)}{\kappa}$
by recalling the formula $\sum_{d\mid\kappa}\Lambda(d)=\log{\kappa}$. Hence, applying (4.1) and (4.3), we find that the estimate
(4.19) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\\ (k_{1},k_{2})>1\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
$\leq\frac{1}{8}g^{m}(\log{y}+O(1))^{m+1}$
holds with the same argument as in (4.18). By (4.18) and (4.19), we derive
$\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
$\leq\frac{1}{8}g^{m}(2g+1)(\log{y}+O(1))^{m+1}.$
One can also obtain
(4.20) $\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
$\leq A(f)^{m+1}(\log{x})^{m+1}\left(\frac{y}{\log{y}}\right)^{m+1}$
in a similar way using (4.4) in place of (4.1) and (4.3). From these, we conclude that the desired estimate for $W_{m}^{+}$ follows. Finally, we have for $2\leq i\leq m$,
$\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{i}\leq h,~{}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{i}}<k_{i+1}\leq f(x)\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{i+1})$
$\leq A(f)^{i+1}(\log{x})^{i}\log{f(x)}\left(\frac{y}{\log{y}}\right)^{i+1}$
along the same line as (4.20). Therefore $W_{i}^{-}$ is estimated as
$W_{i}^{-}\leq\frac{A(f)^{i+1}(\log{x})^{i}\log{f(x)}\left(\frac{y}{\log{y}}\right)^{i+1}}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{i-1}}$
for $2\leq i\leq m$. Furthermore, one can check that the estimate
$W_{1}^{-}\leq\frac{A(f)^{2}(\log{f(x)})^{2}\left(\frac{y}{\log{y}}\right)^{2}}{(\log{f(z)})^{2}}$
holds. Hence the desired estimate for $W_{i}^{-}$ holds for every $i=1,\ldots,m$ by recalling that $\log{f(x)}>\log{f(z)}>\log{x}$. ∎
Proof of Theorem 1.1.
Recall that we may suppose that $f(t)$ satisfies the conditions of Propositions 4.3 and 4.4 without loss of generality. Then we complete the proof by suitably choosing the parameters in Propositions 4.3 and 4.4. Let
(4.21) $y=x^{1/u}\quad\text{and}\quad m=[u]\quad\text{with}\quad 1\leq u\leq\frac{\sqrt{\log{x}}}{\log\log{x}}.$
Then we have
$hg^{m}(\log{y}+O(1))^{m}$ $\leq xg^{[u]}(\log{x})^{[u]}u^{-[u]}\left(1+O\left(\frac{u}{\log{x}}\right)\right)^{[u]}$
$=xg^{[u]}(\log{x})^{[u]}u^{-[u]}\left(1+O\left(\frac{1}{(\log\log{x})^{2}}\right)\right)$
for $x\geq x_{0}(f)$ with a large positive constant $x_{0}(f)$. On the other hand, we obtain
(4.22) $A(f)^{m}(\log{x})^{m}\left(\frac{y}{\log{y}}\right)^{m}$ $\leq xA(f)^{[u]}u^{[u]}$
$=xg^{[u]}(\log{x})^{[u]}u^{-[u]}\left(\frac{A(f)}{g}\frac{u^{2}}{\log{x}}\right)^{[u]}$
$\leq xg^{[u]}(\log{x})^{[u]}u^{-[u]}\frac{A(f)}{(\log\log{x})^{2}}$
for $x\geq\exp(\exp(\sqrt{A(f)}))$. Therefore, Proposition 4.3 implies the upper bound
(4.23) $V_{m}^{+}\leq\frac{xg^{[u]}(\log{x})^{[u]}u^{-[u]}}{2\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}\left(1+O\left(\frac{1}{(\log\log{x})^{2}}\right)\right)$
for sufficiently large $x$, provided the conditions $f(z)>x$ and $x>z>T_{0}(f)$ are satisfied. Then we choose the parameter $z$ as
(4.24) $z=\frac{g^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\frac{x}{\log{x}}.$
One can see from the condition on $u$ in (4.21) that $\sqrt{x}<z<4x(\log{x})^{-1}$ is satisfied for sufficiently large $x$. Therefore the condition $x>z>T_{0}(f)$ holds by making $x$ large enough. To check that $f(z)>x$ holds, we factorize the polynomial $f(t)$ as $f(t)=c_{f}(t-\alpha_{1})\cdots(t-\alpha_{d})$ with $c_{f}\in\mathbb{Z}_{>0}$ and $\alpha_{1},\ldots,\alpha_{d}\in\mathbb{C}$. Then $\log{f(t)}$ is evaluated as
$\log{f(t)}=\log{c_{f}}+d\log{t}+\sum_{j=1}^{d}\log\left|1-\frac{\alpha_{j}}{t}\right|=d\log{t}+O(1)$
for $t>\max\{T_{0}(f),|\alpha_{1}|,\ldots,|\alpha_{d}|\}$, where the implied constant depends only on $c_{f}$ and $d$. Applying this, we derive the asymptotic formula
$\log{f(z)}$ $=d\log{x}+d\log\left(\frac{g^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\right)+O(\log\log{x})$
$=d\log{x}+O\left(u(\log{u}+1)+\log\log{x}\right),$
where the implied constant depends at most on $f(t)$. This yields the formulas
(4.25) $\log{f(z)}$ $=d\log{x}\left(1+O\left(\frac{1}{\sqrt{\log{x}}}\right)\right),$
(4.26) $\log{\frac{f(z)}{x}}$ $=(d-1)\log{x}\left(1+O\left(\frac{1}{\sqrt{\log{x}}}\right)\right)$
by the condition on $u$ in (4.21). Here, we note that (4.25) implies that $f(z)>x$ holds for sufficiently large $x$ due to $d\geq 2$. Furthermore, inserting (4.25) and (4.26) to (4.23), we arrive at
$V_{m}^{+}\leq\frac{xg^{[u]}}{2d(d-1)^{[u]-1}u^{[u]}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right)$
by recalling that we have
$\left(1+O\left(\frac{1}{\sqrt{\log{x}}}\right)\right)^{m}=1+O\left(\frac{1}{\log\log{x}}\right)$
for sufficiently large $x$. Next, we evaluate $V_{i}^{-}$ for $i=1,\ldots,m$ as follows. In a way similar to (4.22), we obtain
$A(f)^{i}(\log{x})^{i-1}\log{f(x)}\left(\frac{y}{\log{y}}\right)^{i}\ll g^{i}x^{i/u}(\log{x})^{i}u^{-i}\frac{1}{(\log\log{x})^{2}},$
where the implied constant depends at most on $f(t)$. Applying Proposition 4.3, we derive the estimate
$V_{i}^{-}\ll\frac{g^{i}x^{i/u}(\log{x})^{i}u^{-i}}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{i-1}}\frac{1}{(\log\log{x})^{2}}\ll\frac{d-1}{d}\left(\frac{g}{d-1}\frac{x^{1/u}}{u}\right)^{i}\frac{1}{(\log\log{x})^{2}}$
by using (4.25) and (4.26), where the implied constant is independent of $i$ and depends at most on $f(t)$. Remark that
$\frac{x^{1/u}}{u}=\exp\left(\frac{1}{u}\log{x}-\log{u}\right)\geq\exp\left(\sqrt{\log{x}}\right)>2$
holds for any $x\geq 2$ by (4.21). It yields
$\sum_{i=1}^{m}V_{i}^{-}$ $\ll\frac{d-1}{d}\left(\frac{g}{d-1}\frac{x^{1/u}}{u}\right)^{m}\frac{1}{(\log\log{x})^{2}}$
$\leq\frac{xg^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\frac{1}{(\log\log{x})^{2}}$
since $A+A^{2}+\cdots+A^{m}<2A^{m}$ for $A>2$. The bounds for $W_{m}^{+}$ and $W_{i}^{-}$ can be obtained similarly. Indeed, one can see that
$W_{m}^{+}$ $\leq\frac{xg^{[u]}(2g+1)}{8d^{2}(d-1)^{[u]-1}u^{[u]+1}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right),$
$\sum_{i=1}^{m}W_{i}^{-}$ $\ll\frac{xg^{[u]}(2g+1)}{d^{2}(d-1)^{[u]-1}u^{[u]+1}}\frac{1}{(\log\log{x})^{2}}$
for sufficiently large $x$. As a result, we conclude that inequality (2.1) holds with
$V$ $=\frac{xg^{[u]}}{2d(d-1)^{[u]-1}u^{[u]}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right),$
$W$ $=\frac{xg^{[u]}(2g+1)}{8d^{2}(d-1)^{[u]-1}u^{[u]+1}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right)$
by Proposition 3.2. Therefore, by (2.2), we obtain
$\Psi_{f}(x,x^{1/u})-\Psi_{f}(z,x^{1/u})$
$<\left(\frac{1}{2}+\frac{1}{2}\frac{2g+1}{8du}+\sqrt{\frac{1}{2}\frac{2g+1}{8du}+\frac{1}{4}\left(\frac{2g+1}{8du}\right)^{2}}\right)$
×$\qquad\times\frac{xg^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right)$
$=\gamma_{f}(u)\frac{g^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(x+O\left(\frac{x}{\log\log{x}}\right)\right).$
We have trivially $\Psi_{f}(z,x^{1/u})\leq z$. Hence the desired bound for $\Psi_{f}(x,x^{1/u})$ follows by the choice of $z$ as in (4.24). ∎
5. Proof of Theorem 1.2
Let $\alpha$ be an arbitrary algebraic irrational number, and $K=\mathbb{Q}(\alpha)$ the algebraic field generated by $\alpha$ over the rationals. Then we define $P_{K}$ as the set of all prime ideals $\mathfrak{p}$ of $K$ such that the following conditions are satisfied:
  • $(\mathrm{i})$
    $\mathfrak{p}$ is of the first degree;
  • $(\mathrm{ii})$
    $\mathfrak{p}$ is unambiguous.
In other words, we have $\operatorname{N}(\mathfrak{p})=p$ and $\mathfrak{p}^{2}\nmid(p)$ for a rational prime $p$ if $\mathfrak{p}\in P_{K}$. Let $\mathfrak{a}$ be the ideal denominator of the fractional ideal $(\alpha)$ of $K$. Then we factorize the ideal $(n+\alpha)\mathfrak{a}$ for $n\in\mathbb{Z}$ as $(n+\alpha)\mathfrak{a}=\mathfrak{a}_{n}\mathfrak{b}_{n}$, where
$\mathfrak{a}_{n}=\prod_{\mathfrak{p}\in P_{K}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}\quad\text{and}\quad\mathfrak{b}_{n}=\prod_{\mathfrak{p}\notin P_{K}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)},$
with $v_{\mathfrak{p}}(n+\alpha)\in\mathbb{Z}_{\geq 0}$ such that $v_{\mathfrak{p}}(n+\alpha)=0$ for all but finitely many of the prime ideals $\mathfrak{p}$. We begin by proving the following preliminary lemmas.
Lemma 5.1.
There exists a constant $C(\alpha)>0$ depending only on $\alpha$ such that
$\operatorname{N}(\mathfrak{b}_{n})\leq C(\alpha)$
holds for any $n\in\mathbb{Z}$.
Proof.
Take a prime ideal $\mathfrak{p}$ of $K$ and $v\geq 1$ such that $\mathfrak{p}^{v}\mid\mathfrak{b}_{n}$. We consider an upper bound for the norm of $\mathfrak{p}^{v}$. Let $L$ denote the Galois closure of $K$ over $\mathbb{Q}$. For a rational prime $p$ with $\mathfrak{p}\mid p$, we write the factorization of $p$ in $L$ as
$p\mathcal{O}_{L}=\left(\mathfrak{P}_{1}\cdots\mathfrak{P}_{g}\right)^{e},$
where $\mathfrak{P}_{1},\ldots,\mathfrak{P}_{g}$ are prime ideals of $L$ with $N(\mathfrak{P}_{k})=p^{f}$. Note that at least one of $e\geq 2$ and $f\geq 2$ is true. Indeed, if $e=f=1$, then $p$ splits completely in $L$. This implies that $p$ splits completely also in $K$, but $\mathfrak{p}\in P_{K}$ follows in this case. Next, we factorize the prime ideal $\mathfrak{p}$ as
(5.1) $\mathfrak{p}\mathcal{O}_{L}=\left(\mathfrak{P}_{k_{1}}\cdots\mathfrak{P}_{k_{m}}\right)^{\eta},$
where $k_{1},\ldots,k_{m}\in\{1,\ldots,g\}$ and $1\leq\eta\leq e$. Denote by $M_{j}$ the decomposition field of $\mathfrak{P}_{k_{j}}$ over $\mathbb{Q}$, and define $\mathfrak{q}_{j}=\mathfrak{P}_{k_{j}}\cap\mathcal{O}_{M_{j}}$. Then $\mathfrak{P}_{k_{j}}$ is the unique prime ideal of $L$ lying over $\mathfrak{q}_{j}$. More precisely, we see that
(5.2) $\mathfrak{q}_{j}\mathcal{O}_{L}=\mathfrak{P}_{k_{j}}^{e}\quad\text{and}\quad\operatorname{N}(\mathfrak{q}_{j})=p$
by the standard properties of decomposition fields. Put $v\eta=eq+r$ with $q\geq 0$ and $0\leq r<e$. From (5.1) and (5.2), we obtain
(5.3) $\mathfrak{p}^{v}\mathcal{O}_{L}$ $=\left(\mathfrak{P}_{k_{1}}\cdots\mathfrak{P}_{k_{m}}\right)^{v\eta}$
$=\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}\cdot\mathfrak{P}_{k_{1}}^{r}\cdots\mathfrak{P}_{k_{m}}^{r}.$
Hence, taking the norms, we have
(5.4) $\operatorname{N}(\mathfrak{p}^{v})^{[L:K]}=\operatorname{N}(\mathfrak{q}_{1}^{q})^{d}\cdots\operatorname{N}(\mathfrak{q}_{m}^{q})^{d}\cdot\operatorname{N}(\mathfrak{P}_{k_{1}})^{r}\cdots\operatorname{N}(\mathfrak{P}_{k_{m}})^{r},$
where $d=[L:M_{j}]=ef$. Denote by $\Delta_{L/\mathbb{Q}}$ the absolute value of the discriminant of $L$ over $\mathbb{Q}$. If $e\geq 2$, then $p\mid\Delta_{L/\mathbb{Q}}$ is satisfied. In this case, $\operatorname{N}(\mathfrak{P}_{k_{j}})^{r}=p^{fr}\leq\Delta_{L/\mathbb{Q}}^{ef}$ holds. If $e=1$, then $\operatorname{N}(\mathfrak{P}_{k_{j}})^{r}=1$ since $r$ must be equal to 00. As a result, we obtain
(5.5) $\operatorname{N}(\mathfrak{P}_{k_{1}})^{r}\cdots\operatorname{N}(\mathfrak{P}_{k_{m}})^{r}\leq\Delta_{L/\mathbb{Q}}^{efm}\leq\Delta_{L/\mathbb{Q}}^{[L:\mathbb{Q}]}$
due to $[L:\mathbb{Q}]=efg$. Next, we choose $\beta\in\mathcal{O}_{K}$ and $c\in\mathbb{Z}_{\geq 1}$ so that $\alpha=\beta/c$. Note that we can determine them uniquely from $\alpha$ by fixing $c$ as the minimum one with this property. Recall that $\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}$ divides the ideal $(n+\alpha)\mathfrak{a}\mathcal{O}_{L}$ by (5.3) due to $\mathfrak{p}^{v}\mid\mathfrak{b}_{n}$. Then we obtain
$\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}\mid(nc+\beta)\mathcal{O}_{L}$
since $(n+\alpha)\mathfrak{a}$ divides $(nc+\beta)$. In other words, $nc+\beta$ belongs to $\mathfrak{q}_{j}^{q}\mathcal{O}_{L}$ for any $j$. Suppose that $\mathcal{O}_{L}=\mathcal{O}_{M_{j}}[x_{1},\ldots,x_{d}]$ with $x_{1},\ldots,x_{d}\in\mathcal{O}_{L}$. Then we have
$1=a_{1}x_{1}+\cdots+a_{d}x_{d}\quad\text{and}\quad\beta=b_{1}x_{1}+\cdots+b_{d}x_{d}$
with some constants $a_{\ell},b_{\ell}\in\mathcal{O}_{M_{j}}$. It yields the formula
$nc+\beta=(na_{1}c+b_{1})x_{1}+\cdots+(na_{d}c+b_{d})x_{d}.$
Then $nc+\beta\in\mathfrak{q}_{j}^{q}\mathcal{O}_{L}$ is equivalent that
$(na_{1}c+b_{1})x_{1}+\cdots+(na_{d}c+b_{d})x_{d}=\sum_{t=1}^{T}p_{t}z_{t}$
with some $p_{t}\in\mathfrak{q}_{j}^{q}$ and $z_{t}\in\mathcal{O}_{L}$. We have
$\sum_{t=1}^{T}p_{t}z_{t}=\sum_{t=1}^{T}p_{t}(A_{t1}x_{1}+\cdots+A_{td}x_{d}),$
where $A_{t1},\ldots,A_{td}\in\mathcal{O}_{M_{j}}$. Comparing them, we derive
$na_{\ell}c+b_{\ell}=\sum_{t=1}^{T}p_{t}A_{t\ell}\in\mathfrak{q}_{j}^{q}$
for any $\ell\in\{1,\ldots,d\}$. This implies that $a_{k}b_{\ell}-a_{\ell}b_{k}\in\mathfrak{q}_{j}^{q}$ for any $k,\ell\in\{1,\ldots,d\}$. Hence we conclude that $\mathfrak{q}_{j}^{q}$ divides the integral ideal generated by all elements in $\{a_{k}b_{\ell}-a_{\ell}b_{k}\mid k\neq\ell\}$ over $\mathcal{O}_{M_{j}}$. Define $G(\alpha;M_{j})$ as the norm of this ideal. One can check that $G(\alpha;M_{j})$ is independent to the choice of the basis $\{x_{1},\ldots,x_{d}\}$. Furthermore, $G(\alpha)=\max G(\alpha;M)$ depends only on $\alpha$, where the maximum is taken over all algebraic fields $M$ such that $\mathbb{Q}\subset M\subsetneq L$. Finally, we see that
(5.6) $\operatorname{N}(\mathfrak{q}_{1}^{q})^{d}\cdots\operatorname{N}(\mathfrak{q}_{m}^{q})^{d}\leq G(\alpha)^{dm}\leq G(\alpha)^{[L:\mathbb{Q}]}$
due to $d=ef$ and $[L:\mathbb{Q}]=efg$. By (5.4), (5.5), and (5.6), we arrive at
$\operatorname{N}(\mathfrak{p}^{v})\leq\left(G(\alpha)\Delta_{L/\mathbb{Q}}\right)^{[K:\mathbb{Q}]}=C_{1}(\alpha),$
say. Then the desired result immediately follows. Indeed, we have
$\operatorname{N}(\mathfrak{b}_{n})=\prod_{\mathfrak{p}\mid\mathfrak{b}_{n}}\operatorname{N}(\mathfrak{p}^{\nu_{\mathfrak{p}}(\mathfrak{b}_{n})})\leq C_{1}(\alpha)^{\pi_{K}(C_{1}(\alpha))},$
where $\pi_{K}(x)$ denotes the number of prime ideals $\mathfrak{p}$ of $K$ such that $\operatorname{N}(\mathfrak{p})\leq x$. ∎
Lemma 5.2.
Take an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$, and denote by $A_{f}$ its leading coefficient. Let $C(\alpha)>0$ be the constant described in Lemma 5.1. Then, for any $n\in\mathbb{Z}$ and any $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$, the following conditions are equivalent.
  • $(\mathrm{A})$
    There exists a rational prime $p$ such that $p\mid f(n)$ and $p>x$.
  • $(\mathrm{B})$
    There exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
Proof.
We first remark that the identity
(5.7) $\frac{\operatorname{N}((n+\alpha)\mathfrak{a})}{\operatorname{N}(\mathfrak{a})}=\left|(n+\alpha_{1})\cdots(n+\alpha_{d})\right|=\left|\frac{f(n)}{A_{f}}\right|$
holds, where $\alpha_{1},\ldots,\alpha_{d}$ are the conjugates of $\alpha$ over $\mathbb{Q}$. Suppose that there exists a rational prime $p$ such that $p\mid f(n)$ and $p>x$. By (5.7), we deduce that $p$ divides $\operatorname{N}((n+\alpha)\mathfrak{a})$ or $A_{f}$, but the later does not hold due to $x>|A_{f}|$. Thus we have that $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$. Furthermore, it implies that there exists a prime ideal $\mathfrak{p}$ of $K$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})=p^{f}$ with some $j\geq 1$. If we suppose that $\mathfrak{p}\notin P_{K}$, then $\operatorname{N}(\mathfrak{p})>x>C(\alpha)$ would follow, which yields contradiction by Lemma 5.1. Hence we conclude that $\mathfrak{p}\in P_{K}$ holds, and therefore, $(\mathrm{A})$ implies $(\mathrm{B})$.
Suppose conversely that there exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$. Then $\operatorname{N}(\mathfrak{p})=p$ holds for a rational prime $p$ by the definition of $P_{K}$. Thus $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$ and $p>x$ are satisfied. Furthermore, $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$ implies that $p$ divides $f(n)$ or $\operatorname{N}(\mathfrak{a})$ by (5.7), but the later does not hold due to $x>\operatorname{N}(\mathfrak{a})$. As a result, we conclude that $(\mathrm{B})$ implies $(\mathrm{A})$. ∎
Lemma 5.3.
Let $\mathfrak{p}\in P_{K}$ and $v\geq 1$. Assume $\mathfrak{p}^{v}\mid(m+\alpha)\mathfrak{a}$ and $\mathfrak{p}^{v}\mid(n+\alpha)\mathfrak{a}$ for $m,n\in\mathbb{Z}$. Then we have $m\equiv n~{}(\bmod\,{p^{v}})$ with $\operatorname{N}(\mathfrak{p})=p$.
Proof.
Put $m-n=p^{w}A$ with some $w\in\mathbb{Z}_{\geq 0}$ and $A\in\mathbb{Z}_{\geq 1}$ such that $(p,A)=1$. By the assumption, we have $(m+\alpha)\mathfrak{a}\subset\mathfrak{p}^{v}$ and $(n+\alpha)\mathfrak{a}\subset\mathfrak{p}^{v}$. Thus we see that $(m-n)\mathfrak{a}\subset\mathfrak{p}^{v}$ holds, which yields
$\mathfrak{p}^{v}\mid(m-n)$
due to $\mathfrak{p}\nmid\mathfrak{a}$. Therefore, $\mathfrak{p}^{v}\mid(p)^{w}(A)$ follows. We deduce that $\mathfrak{p}\nmid(A)$ from $(p,A)=1$ by taking the norms. As a result, we obtain
$\mathfrak{p}^{v}\mid(p)^{w}.$
Recall that $\mathfrak{p}$ is unambiguous, i.e. $\mathfrak{p}^{2}\nmid(p)$. Hence we see that $w\geq v$, which implies $p^{v}\mid(m-n)$ as desired. ∎
From the above, we associate $C_{\alpha}(x)$ defined in Section 1.2 with $\Psi_{f}(x,y)$ for an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$. The relation between $C_{\alpha}(x)$ and $\Psi_{f}(x,y)$ was mentioned firstly by Worley [20], and a more detailed research was presented by Andersson in his unpublished work [1]. In this paper, we obtain the following result.
Proposition 5.4.
Take an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$. Then the asymptotic formula
$C_{\alpha}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x}{\log{x}}\right)$
holds for sufficiently large $x$, where the implied constant depends only on $K=\mathbb{Q}(\alpha)$.
Proof.
Denote by $M_{\alpha}(x)$ the set of integers $n$ in $[1,x]$ such that there exists a prime ideal $\mathfrak{p}$ of $K$ with conditions (1.5). Furthermore, we also define $N_{f}(x)$ as the set of integers $n$ in $[1,x]$ such that $f(n)$ is not $x$-smooth. Then we have
(5.8) $|M_{\alpha}(x)|=C_{\alpha}(x)\quad\text{and}\quad|N_{f}(x)|=x-\Psi_{f}(x,x)+O(1)$
with an absolute implied constant. Take an integer $n\in N_{f}(x)$ arbitrarily. Then there exists a rational prime $p$ such that $p\mid f(n)$ and $p>x$. Here, we suppose $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$ as in Lemma 5.2. Then we deduce from Lemma 5.2 that there exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$. If $\mathfrak{p}\mid(m+\alpha)\mathfrak{a}$ is satisfied for some $m\in\mathbb{Z}$ in $[1,x]$, then we apply Lemma 5.3 to deduce that $m\equiv n~{}(\bmod\,{p})$ with $\operatorname{N}(\mathfrak{p})=p$. Since $m,n\in[1,x]$ and $p>x$, we must obtain $m=n$, and therefore, the condition
$\mathfrak{p}\nmid\prod_{\begin{subarray}{c}1\leq m\leq x\\ m\neq n\end{subarray}}(m+\alpha)\mathfrak{a}$
is satisfied. As a result, we find that $n\in M_{\alpha}(x)$ holds if $n\in N_{f}(x)$. Hence the inclusion $N_{f}(x)\subset M_{\alpha}(x)$ follows.
Next, we define $\phi(n)$ for $n\in\mathbb{Z}$ as the set of all prime ideals $\mathfrak{p}$ of $K$ such that conditions (1.5) are satisfied. Then it is non-empty for any $n\in M_{\alpha}(x)$ by definition. On the other hand, we have $\phi(m)\cap\phi(n)=\emptyset$ for any integers $m,n\in M_{\alpha}(x)$ with $m\neq n$ by conditions (1.5). Furthermore, if $n\in M_{\alpha}(x)\setminus N_{f}(x)$, then any prime ideal $\mathfrak{p}\in\phi(n)$ satisfies $\operatorname{N}(\mathfrak{p})\leq x$ by Lemma 5.2. Hence, we see that
$\left|M_{\alpha}(x)\setminus N_{f}(x)\right|\leq\pi_{K}(x)\ll\frac{x}{\log{x}}$
by the prime ideal theorem, where the implied constant depends only on $K$. It yields the desired result by formulas (5.8). ∎
Proof of Theorem 1.2.
Let $f(t)\in\mathbb{Z}[t]$ be an irreducible polynomial with $f(-\alpha)=0$. Then $f(t)$ is of degree $d=\deg(\alpha)\geq 2$. By Proposition 5.4 and Theorem 1.1 with $u=1$, we have
$C_{\alpha}(x)>x-\frac{\gamma_{f}(1)}{d}\left(x+O\left(\frac{x}{\log\log{x}}\right)\right)+O\left(\frac{x}{\log{x}}\right),$
where the implied constants depend only on $f(t)$ and $K$. It yields the inequality
$C_{\alpha}(x)\geq\left(1-\frac{\gamma_{f}(1)}{d}\right)x$
for sufficiently large $x$. Furthermore, the coefficient is calculated as
$1-\frac{\gamma_{f}(1)}{d}=1-\frac{1}{2d}-\frac{3}{16d^{2}}-\frac{1}{d}\sqrt{\frac{3}{16d}+\left(\frac{3}{16d}\right)^{2}}$
by using (1.3). We also obtain
$0.543\displaystyle 1-\frac{\gamma_{f}(1)}{d}>1-\frac{0.914}{2}=$ 0.543
by (1.4) and $d\geq 2$. Hence the proof is completed. ∎
Let $f(t)$ be an irreducible polynomial with $f(-\alpha)=0$. If (1.1) is true at least for $u=1$, then we deduce from Proposition 5.4 that
$C_{\alpha}(x)\sim(1-\rho(d))x$
as $x\to\infty$, where we put $d=\deg(\alpha)\geq 2$. Here, it is known that $\rho(d)=d^{-d+o(d)}$ as $d\to\infty$. Thus the result of Theorem 1.2 still leaves room for improvement.
Remark 5.5.
Actually, Lemma 5.1 is much stronger than necessary for the proof of Theorem 1.2 since one needs just that $\operatorname{N}(\mathfrak{p})$ is bounded for any $\mathfrak{p}\notin P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$. We note that Lemma 5.1 is a modification of a statement of Cassels [2]. He used the set $P_{K}^{*}$ consisting of prime ideals $\mathfrak{p}$ of $K$ such that
  • $(\mathrm{i})$
    $\mathfrak{p}$ is of the first degree;
  • $(\mathrm{ii})$
    $\mathfrak{p}$ is unambiguous;
  • $(\mathrm{iii})$
    for any $m\in\mathbb{Z}$, if $\mathfrak{p}\mid(m+\alpha)\mathfrak{a}$ then $(m+\alpha)\mathfrak{a}$ is not divisible by any prime ideal $\mathfrak{p}^{\prime}\neq\mathfrak{p}$ satisfying $\operatorname{N}(\mathfrak{p})=\operatorname{N}(\mathfrak{p}^{\prime})$.
We factorize the ideal $(n+\alpha)\mathfrak{a}$ as $(n+\alpha)\mathfrak{a}=\mathfrak{a}_{n}^{*}\mathfrak{b}_{n}^{*}$, where
$\mathfrak{a}_{n}^{*}=\prod_{\mathfrak{p}\in P_{K}^{*}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}\quad\text{and}\quad\mathfrak{b}_{n}^{*}=\prod_{\mathfrak{p}\notin P_{K}^{*}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}.$
Then it was claimed in [2, p. 181] that $\operatorname{N}(\mathfrak{b}_{n}^{*})$ is bounded for $n\in\mathbb{Z}$. However, we know that $\operatorname{N}(\mathfrak{b}_{n}^{*})$ is not bounded in general. For example, letting $\alpha=7\sqrt{2}$, one can find for any $v\geq 1$ an integer $n=n(v)$ such that $\operatorname{N}(b_{n}^{*})\geq 7^{v}$. See [16, Remark 1.6] for more information. Hence the original method by Cassels seems flawed.
The boundedness of $\operatorname{N}(\mathfrak{b}_{n}^{*})$ was used in [2] to deduce the inequality
$\sum_{\mathfrak{p}\in P_{K}^{*}}v_{\mathfrak{p}}(n+\alpha)\log{\operatorname{N}(\mathfrak{p})}>2\log{n}-C$
for a constant $C>0$ which is independent of $n$. Furthermore, condition $(\mathrm{iii})$ was used to see that there is no contribution of terms for $\mathfrak{p}_{1},\mathfrak{p}_{2}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})$ in the sum
(5.9) $\sum_{n\in\mathfrak{S}_{N}}\sum_{\begin{subarray}{c}\mathfrak{p}_{1},\mathfrak{p}_{2}\in P_{K}^{*}\\ \mathfrak{p}_{j}\mid(n+\alpha)\mathfrak{a}\\ \operatorname{N}(\mathfrak{p}_{j})<\sqrt{M}\end{subarray}}\frac{\log{\operatorname{N}(\mathfrak{p}_{1})}}{\operatorname{N}(\mathfrak{p}_{1})}\frac{\log{\operatorname{N}(\mathfrak{p}_{2})}}{\operatorname{N}(\mathfrak{p}_{2})},$
where $\mathfrak{S}_{N}$ is the set of integers $n$ in $[N,N+M]$ such that $\operatorname{N}(\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)})<M$ holds for all $\mathfrak{p}\in P_{K}^{*}$ with $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$. Note that (5.9) appeared during a calculation using the Cauchy–Schwarz inequality, and that a similar argument was also presented in the proof of Proposition 2.1. The difference is that the method of this paper is non ideal-theoretic, which requires nothing similar to condition $(\mathrm{iii})$. On the other hand, it is also possible to recover the proof in [2] by replacing $P_{K}^{*}$ with $P_{K}$. Indeed, $\operatorname{N}(\mathfrak{b}_{n})$ is bounded by Lemma 5.1, which yields
$\sum_{\mathfrak{p}\in P_{K}}v_{\mathfrak{p}}(n+\alpha)\log{\operatorname{N}(\mathfrak{p})}>2\log{n}-C.$
Furthermore, for any prime ideals $\mathfrak{p}_{1},\mathfrak{p}_{2}\in P_{K}$ with $\mathfrak{p}_{j}\mid(n+\alpha)\mathfrak{a}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})=:p$, one can show that $p$ is bounded for $n\in\mathbb{Z}$. Actually, this is sufficient to show that the contribution of terms for $\mathfrak{p}_{1},\mathfrak{p}_{2}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})$ in (5.9) is not zero but can be negligible.
To close this section, we confirm the original statement of Cassels [2].
Theorem 5.6 (Cassels [2]).
Let $\alpha$ be an algebraic irrational number. Then there exists an integer $N_{0}=N_{0}(\alpha)>10^{6}$ depending only on $\alpha$ with the following property. Suppose that $N\geq N_{0}$ and put $M=\lfloor 10^{-6}N\rfloor$. Then at least $10051M/$ 100 integers in $N<n\leq N+M$ are such that $(n+\alpha)\mathfrak{a}$ is divisible by a prime ideal $\mathfrak{p}$ which does not divide $(m+\alpha)\mathfrak{a}$ for any integer $0\leq m\leq N+M$ with $m\neq n$.
Proof.
The only difference is the intervals where $n$ belongs. To show Theorem 5.6, we apply Propositions 4.3 and 4.4 with $x=N+M$, $z=M$, and $m=u=1$. Then we see that inequality (2.1) holds with
$V$ $=\frac{M(\log(N+M)+O(1))}{2\log{f(N)}}+O\left(\frac{N+M}{\log{N}}\right),$
$W$ $=\frac{3M(\log(N+M)+O(1))^{2}}{8(\log{f(N)})^{2}}+O\left(\frac{N+M}{(\log{N})^{2}}\right)$
for any polynomial $f(t)$ satisfying the assumptions of Proposition 4.3 and 4.4. Thus we obtain
$\Psi_{f}(N+M,N+M)-\Psi_{f}(N,N+M)<\frac{\kappa_{f}(1)}{d}M<0.48M$
by (2.2) if $N$ is large enough. Furthermore, along the same line as the proof of Proposition 5.4, we see that the number of integers in $N<n\leq N+M$ with the desired property is equal to
$M-\left(\Psi_{f}(N+M,N+M)-\Psi_{f}(N,N+M)\right)+O\left(\frac{N+M}{\log(N+M)}\right),$
where $f(t)\in\mathbb{Z}[t]$ is an irreducible polynomial with $f(-\alpha)=0$. Hence we obtain the conclusion. ∎
6. Proof of Theorems 1.3 and 1.4
We prove Theorems 1.3 and 1.4 by associating the functions $R_{b}(x)$ and $N(x)$ defined in Section 1.3 with $\Psi_{f}(x,y)$ for a suitable polynomial $f(t)$. For this, we recall the following criterions.
Lemma 6.1 (Everest–Stevens–Tamsett–Ward [7]).
Let $n\in\mathbb{Z}_{>|b|}$ with $-b$ not an integer square. Then $n^{2}+b$ has a primitive divisor if and only if $P^{+}(n^{2}+b)>2n$.
Lemma 6.2 (Todd [19]).
Let $n\in\mathbb{Z}_{\geq 1}$. Then $\arctan n$ is irreducible if and only if $P^{+}(n^{2}+1)>2n$.
Then the equality $N(x)=R_{1}(x)$ immediately follows. Furthermore, we associate $R_{b}(x)$ with $\Psi_{f}(x,y)$ for the quadratic polynomial $f(t)=t^{2}+b$ as follows.
Proposition 6.3.
Let $f(t)=t^{2}+b$ with $-b$ not an integer square. Then we have
$R_{b}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x\log\log{x}}{\log{x}}\right)$
for sufficiently large $x$, where the implied constant depends only on $|b|$.
Proof.
By Lemma 6.1, $R_{b}(x)$ is evaluated as
$R_{b}(x)=x-\sum_{\begin{subarray}{c}1\leq n\leq x\\ P^{+}(n^{2}+b)<2n\end{subarray}}1+O(1)$
for $x\geq 1$, where the implied constant depends only on $|b|$. Furthermore, we obtain
$\sum_{\begin{subarray}{c}1\leq n\leq x\\ P^{+}(n^{2}+b)<2n\end{subarray}}1$ $=\sum_{\begin{subarray}{c}1\leq n\leq x/2\\ P^{+}(n^{2}+b)<2n\end{subarray}}1+\sum_{\begin{subarray}{c}x/2<n\leq x\\ P^{+}(n^{2}+b)<2n\end{subarray}}1$
$=\sum_{\begin{subarray}{c}1\leq n\leq x/2\\ P^{+}(n^{2}+b)\leq x\end{subarray}}1-\sum_{\begin{subarray}{c}1\leq n\leq x/2\\ 2n\leq P^{+}(n^{2}+b)\leq x\end{subarray}}1+\sum_{\begin{subarray}{c}x/2<n\leq x\\ P^{+}(n^{2}+b)\leq x\end{subarray}}1+\sum_{\begin{subarray}{c}x/2<n\leq x\\ x<P^{+}(n^{2}+b)<2n\end{subarray}}1$
$=S_{1}-S_{2}+S_{3}+S_{4},$
say. Then $S_{1}+S_{3}=\Psi_{f}(x,x)$ holds with $f(t)=t^{2}+1$ by definition. On the other hand, $S_{2}$ and $S_{4}$ are evaluated as
$S_{2}$ $\leq\sum_{\begin{subarray}{c}1\leq n\leq z\\ 2n\leq P^{+}(n^{2}+b)\leq x\end{subarray}}1+\sum_{\begin{subarray}{c}z\leq n\leq x/2\\ 2n\leq P^{+}(n^{2}+b)\leq x\end{subarray}}1\leq z+\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)\leq x\end{subarray}}1,$
$S_{4}$ $\leq\sum_{\begin{subarray}{c}z<n\leq x\\ 2z<P^{+}(n^{2}+b)<2x\end{subarray}}1$
with a parameter $z$ such that $21\leq z\leq x/$ 2. From the above, we obtain the formula
(6.1) $R_{b}(x)=x-\Psi_{f}(x,x)+O\left(z+\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1\right)$
for $x\geq 1$. Here, we see that
$\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1=\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ P^{+}(n^{2}+b)=p\end{subarray}}1\leq\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0~{}\pmod{p}\end{subarray}}1$
holds since $P^{+}(n^{2}+b)=p$ implies $f(n)\equiv 0~{}(\bmod\,{p})$. We have $\omega_{f}(p)\leq 4\sqrt{|b|}$ by applying (4.5). Therefore, the last sum is further evaluated as
$\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0\pmod{p}\end{subarray}}1$ $\ll\sum_{2z\leq p<2x}\left(\frac{x}{p}+1\right)$
$=x\log\left(\frac{\log(2x)}{\log(2z)}\right)+O\left(\frac{x}{\log(2z)}\right)$
by using the well-known formula
$\sum_{p\leq t}\frac{1}{p}=\log\log{t}+A+O\left(\frac{1}{\log{t}}\right).$
Here, we choose the parameter $z$ as $z=x/(2\log{x})$. In this case, we have
$\log\left(\frac{\log(2x)}{\log(2z)}\right)=\log\left\{\left(1+\frac{\log{2}}{\log{x}}\right)\left(1-\frac{\log\log{x}}{\log{x}}\right)^{-1}\right\}\ll\frac{\log\log{x}}{\log{x}}$
for sufficiently large $x$. Hence we arrive at the estimate
$\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1\ll\frac{x\log\log{x}}{\log{x}},$
which yields the desired result by (6.1). ∎
If (1.1) is true for $f(t)=t^{2}+b$ and $u=1$, then Proposition 6.3 yields that
$R_{b}(x)\sim(1-\rho(2))x$
as $x\to\infty$. Recall that $\rho(u)=1-\log{u}$ for $1\leq u\leq 2$. Hence (1.1) contains the conjectures of Everest–Harman and Chowla–Todd on $R_{b}(x)$ and $N(x)$, respectively.
Proof of Theorems 1.3 and 1.4.
The proof is similar to that of Theorem 1.2. Let $f(t)=t^{2}+b$. By Proposition 6.3 and Theorem 1.1 with $u=1$, we have
$R_{b}(x)>x-\frac{\gamma_{f}(1)}{2}\left(x+O\left(\frac{x}{\log\log{x}}\right)\right)+O\left(\frac{x}{\log{x}}\right),$
which implies the inequality
$R_{b}(x)\geq\left(1-\frac{\gamma_{f}(1)}{2}\right)x$
for sufficiently large $x$. Since we have
$0.543\displaystyle 1-\frac{\gamma_{f}(1)}{2}>1-\frac{0.914}{2}=$ 0.543
by using (1.4), we obtain the result for $R_{b}(x)$. The result for $N(x)$ also follows due to $N(x)=R_{1}(x)$ as explained before. ∎

### v2

@@
 Abstract.
-We provide a new upper bound for the number of smooth values of a polynomial with integer coefficients.
+We prove a new upper bound for the number of smooth values of a polynomial with integer coefficients.
 This improves Timofeev’s previous result unless the polynomial is a product of linear polynomials with integer coefficients.
-As an application, we fix an error in the proof of a result of Cassels which was used to prove that the Hurwitz zeta-function with algebraic irrational parameter has infinitely many zeros on the domain of convergence.
+As an application, we provide another proof for a result of Cassels which was used to prove that the Hurwitz zeta-function with algebraic irrational parameter has infinitely many zeros on the domain of convergence.
 We also apply the main result to a problem on primitive divisors of quadratic polynomials.
@@
 holds in the range $e\leq u\leq(\log{x})(\log\log{x})^{-1}$ for sufficiently large $x$, where $c(f)$ is a positive constant depending on the degree and coefficients of $f(t)$.
+See also a recent article of Suzuki [18] for an improvement of Hmyrova’s estimate.
 For more general polynomials, Timofeev [19] obtained that
@@
 (1.4) $\gamma_{f}(u)=\frac{1}{2}+\frac{3}{16du}+\sqrt{\frac{3}{16du}+\left(\frac{3}{16du}\right)^{2}}\leq\frac{19+\sqrt{105}}{32}=0.913$...
-holds for any $u\geq 1$. This is also useful in applications of Theorem 1.1.
+holds for any $u\geq 1$. While this may seem like only a slight improvement over the previous result, we will see in Sections 1.2 and 1.3 that it has valuable applications to several issues.
 Theorem 1.1 is proved by induction on $m=[u]$.
@@
 (1.5) $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}\quad\text{and}\quad\mathfrak{p}\nmid\prod_{\begin{subarray}{c}1\leq m\leq x\\ m\neq n\end{subarray}}(m+\alpha)\mathfrak{a}.$
-In Section 5, we prove the following result by applying Theorem 1.1.
+A lower bound for $C_{\alpha}(x)$ was used to study the zeros of the Hurwitz zeta-function $\zeta(s,\alpha)$ for $\operatorname{Re}(s)>1$, where $\alpha$ is an algebraic irrational number with $0<\alpha<1$.
+Indeed, Cassels [2] applied the estimate $C_{\alpha}(x)>0.51x$ for large $x$ to showed that $\zeta(s,\alpha)$ has infinitely many zeros in the strip $1<\operatorname{Re}(s)<1+\delta$ for any $\delta>0$, which extends the classical result of Davenport and Heilbronn [4].
+The relation between $C_{\alpha}(x)$ and the number of smooth values of a polynomial $f(t)\in\mathbb{Z}[t]$ was firstly mentioned by Worley [21], and a more detailed research was presented by Andersson in his unpublished work [1].
+In Section 5, we prove the following result.
 Theorem 1.2.
+Take an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$. Then the asymptotic formula
+$C_{\alpha}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x}{\log{x}}\right)$
+holds for sufficiently large $x$, where the implied constant depends only on $K=\mathbb{Q}(\alpha)$.
+Therefore, to show a lower (resp.
+upper) bound for $C_{\alpha}(x)$ is nothing other than to show an lower (resp.
+upper) bound for $\Psi_{f}(x,x)$.
+In particular, Theorem 1.1 implies the following result.
+Corollary 1.3.
 Let $\alpha$ be an arbitrary algebraic irrational number. Then we have
@@
 holds for sufficiently large $x$.
-Theorem 1.2 was essentially obtained by Cassels [2].
-He applied this kind of result to study the zeros of the Hurwitz zeta-function $\zeta(s,\alpha)$ for $\operatorname{Re}(s)>1$, where $\alpha$ is an algebraic irrational number with $10<\alpha<$ 1.
-Indeed, he showed that $\zeta(s,\alpha)$ has infinitely many zeros in the strip $1<\operatorname{Re}(s)<1+\delta$ for any $\delta>0$, which extends the classical result of Davenport and Heilbronn [4].
-However, the proof in [2] contains an error which cannot be easily fixed as explained in [16].
-See Remark 5.5 for more details of the difference between the methods of [2] and this paper.
-Remark that the formulation of Theorem 1.2 is due to Worley [20], and the original statement by Cassels in [2] is slightly different. We discuss how Cassels’ result is proved at the end of Section 5.
+Corollary 1.3 was essentially obtained by Cassels [2] and Worley [21].
+Note that the original proofs in [2, 21] should be modified due to an erroneous classification of prime ideals.
+See [16, Remark 4.10] for the details.
+The main difference between the methods in [2, 21] and this paper is that the proof of Theorem 1.1 does not require any ideal-theoretic arguments.
 1.3. Primitive divisors of $n^{2}+b$
@@
 We say that an integer $d>1$ is a primitive divisor of $A_{n}$ if $d\mid A_{n}$ and $(d,A_{m})=1$ for all nonzero terms $A_{m}$ with $m<n$.
-Then we define 111 In the previous papers, $R_{b}(x)$ was denoted by $\rho_{b}(x)$, and usually $\rho_{1}(x)$ was abbreviated to $\rho(x)$.
+Then we define  [^1] In the previous papers, $R_{b}(x)$ was denoted by $\rho_{b}(x)$, and usually $\rho_{1}(x)$ was abbreviated to $\rho(x)$.
 We change the symbols so as to keep that $\rho$ represents the Dickman function.
 $R_{b}(x)=\#\left\{n\in\mathbb{Z}\cap[1,x]~\middle|~\textrm{$n^{2}+b$ has a primitive divisor}\right\}$
-for $b\in\mathbb{Z}$. Suppose that $-b$ is not an integer square. Then Everest and Harman [6] conjectured that $R_{b}(x)\sim(\log{2})x$ as $x\to\infty$. They also proved
+for $b\in\mathbb{Z}$. Suppose that b-b is not an integer square. Then Everest and Harman [6] conjectured that $R_{b}(x)\sim(\log{2})x$ as $x\to\infty$. They also proved
 $0.5324x<R_{b}(x)<0.905x$
@@
 $0.5377x<R_{1}(x)<0.86x.$
-Li [14] also announced an even better upper bound $R_{1}(x)<0.8437x$. In this paper, we provide a new lower bound for $R_{b}(x)$.
+Li [14] also announced an even better upper bound $R_{1}(x)<0.8437x$. In Section 6, we connect $R_{b}(x)$ with the number of smooth values of the quadratic polynomial $f(t)=t^{2}+b$.
 Theorem 1.4.
-Suppose that $-b$ is not an integer square. Then the inequality
+Let $f(t)=t^{2}+b$ with b-b not an integer square. Then we have
+$R_{b}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x\log\log{x}}{\log{x}}\right)$
+for sufficiently large $x$, where the implied constant depends only on $|b|$.
+Applying Theorem 1.1, we prove a new lower bound for $R_{b}(x)$.
+Corollary 1.5.
+Suppose that b-b is not an integer square. Then the inequality
 $R_{b}(x)>0.543x$
 holds for sufficiently large $x$.
-The proof of Theorem 1.3 is presented in Section 6.
-It should be noted that the method for the proof is applicable only to the lower bound for $R_{b}(x)$, and we have no contribution to the upper bound.
-For $b=1$, Theorem 1.3 is also related to a problem on arctangent relations which was originally studied by Todd [19].
+For $b=1$, Corollary 1.5 is also related to a problem on arctangent relations which was originally studied by Todd [20].
 For $n\in\mathbb{Z}_{\geq 1}$, we say that $\arctan n$ is reducible if there exist integers $f_{1},\ldots,f_{k}$ and $n_{1},\ldots,n_{k}$ with $1\leq n_{j}<n$ such that
@@
 as $x\to\infty$ for any $\epsilon>0$. Actually, it is not difficult to see that $N(x)=R_{1}(x)$ holds; see Lemmas 6.1 and 6.2. Hence we obtain the following result.
-Theorem 1.4.
+Corollary 1.6.
 For sufficiently large $x$, we have $N(x)>0.543x$.
-It would be worth noting that $N(x)<0.8437x$ also follows from Li’s work [14].
+It would be worth noting that $N(x)<0.8437x$ also follows from Li’s work [14]. If (1.1) is true for $f(t)=t^{2}+b$ and $u=1$, then Theorem 1.4 yields that
+$R_{b}(x)\sim(1-\rho(2))x$
+as $x\to\infty$. Recall that $\rho(u)=1-\log{u}$ for $1\leq u\leq 2$. Hence (1.1) contains the conjectures of Everest–Harman and Chowla–Todd on $R_{b}(x)$ and $N(x)$, respectively.
 Acknowledgments
@@
 The author would also like to thank Hiroshi Mikawa.
-Theorems 1.3 and 1.4 were motivated by his lecture at Hachiouji Number Theory Seminar 2024.
+Corollaries 1.5 and 1.6 were motivated by his lecture at Hachiouji Number Theory Seminar 2024.
 The work of this paper was supported by JSPS Grant-in-Aid for Early-Career Scientists (Grant Number 24K16906).
@@
 $W$ $=\frac{1}{(\log{f(z)})^{2}}\sum_{\begin{subarray}{c}1\leq k_{1},k_{2}\leq f(x)\\ P^{+}(k_{1})\leq\sqrt{y}\\ P^{+}(k_{2})\leq\sqrt{y}\end{subarray}}\Lambda(k_{1})\Lambda(k_{2})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1,$
-where $[k_{1},k_{2}]$ denotes the largest common multiple of $k_{1}$ and $k_{2}$.
+where $k_{1},k_{2}]$ denotes the least common multiple of $k_{1}$ and $k_{2}$.
 Proof.
@@
 Recall that $f(j\kappa+u)\equiv f(u)~(\bmod\,{\kappa})$ for any $j\in\mathbb{Z}$. Therefore we obtain
-(3.2) $S(\kappa)\leq\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\end{subarray}}1=\sum_{\begin{subarray}{c}u\pmod{\kappa}\\ f(u)\equiv 0\pmod{\kappa}\end{subarray}}\sum_{\begin{subarray}{c}z<n\leq x\\ n\equiv u\pmod{\kappa}\end{subarray}}1\leq\omega_{f}(\kappa)\left(\frac{x-z}{\kappa}+1\right),$
+(3.2) $S(\kappa)\leq\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\end{subarray}}1=\sum_{\begin{subarray}{c}u\hskip-5.69054pt\pmod{\kappa}\\ f(u)\equiv 0\hskip-5.69054pt\pmod{\kappa}\end{subarray}}\sum_{\begin{subarray}{c}z<n\leq x\\ n\equiv u\hskip-5.69054pt\pmod{\kappa}\end{subarray}}1\leq\omega_{f}(\kappa)\left(\frac{x-z}{\kappa}+1\right),$
 but this estimate is not good when $y$ is small since we have ignored the condition $P^{+}(f(n))\leq y$ in the first inequality.
@@
 Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$.
-Suppose that $f(z)>x$ is satisfied, and put $h=x-z$.
+Suppose that $f(z)>x$ is satisfied, and put h=xzh=x-z.
 Then we have
@@
 Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$.
-Suppose that $f(z)>x$ is satisfied, and put $h=x-z$.
+Suppose that $f(z)>x$ is satisfied, and put h=xzh=x-z.
 Then inequality (2.1) holds with
@@
 $W_{m}^{+}$ $=\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}$
-×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
+$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~(j=1,2)\\ P^{+}(k_{j})\leq y~(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
 with the convention $k_{3}\cdots k_{m+1}=1$ for $m=1$. Secondly, we define $V_{i}^{-}$ as
@@
 $W_{i}^{-}$ $=\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{i-1}}$
-×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{i}\leq h,~{}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{i}}<k_{i+1}\leq f(x)\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{i+1})$
+$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{i}\leq h,~\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{i}}<k_{i+1}\leq f(x)\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{i+1})$
 for every $2\leq i\leq m$ with the convention $k_{3}\cdots k_{i}=1$ for $i=2$, and
@@
 $V_{m}^{+}$ $<\frac{1}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
-×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{k_{1}\cdots k_{m}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ k_{1}\cdots k_{m}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
+$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~(j\geq 2)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{k_{1}\cdots k_{m}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ k_{1}\cdots k_{m}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
 $\quad+\frac{1}{\log{f(z)}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
-×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}\frac{h}{k_{1}\cdots k_{m}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
+$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}\frac{h}{k_{1}\cdots k_{m}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
 Replacing $\lambda$ by $k_{m+1}$, we see that the first and second terms of the right-hand side are equal to $V_{m+1}^{+}$ and $V_{m+1}^{-}$, respectively.
@@
 $W_{m}^{+}<\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
-×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
+$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~(j=1,2)\\ P^{+}(k_{j})\leq y~(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
 $\quad+\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m}}$
-×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
+$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
 In this case, we replace $\lambda$ with $k_{m+2}$.
@@
 Hence we do not expect that $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds unless $k_{1},\ldots,k_{m}$ are pairwise coprime.
-222 It seems to be misunderstood often in [18] that the identity $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds for any $k_{1},\ldots,k_{m}\in\mathbb{Z}_{\geq 1}$.
+[^2] It seems to be misunderstood often in [19] that the identity $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds for any $k_{1},\ldots,k_{m}\in\mathbb{Z}_{\geq 1}$.
 One can fix the errors with the same argument presented in this paper using Lemma 4.2.
@@
 Suppose further that $f(t)\to\infty$ as $t\to\infty$, and that $f(z)>x$.
-Put $h=x-z$.
+Put h=xzh=x-z.
 Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
@@
 $\leq\sum_{\begin{subarray}{c}1\leq k\leq x\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\left(\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\right)^{n-1}\left(d\sqrt{\Delta_{f}}\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\\ (k,\Delta_{f})>1\end{subarray}}\frac{\Lambda(k)}{k}\right)^{m-n}$
-by recalling that $h=x-z<x$. Applying (4.1) and (4.2) of Lemma 4.1, we have
+by recalling that h=xz<xh=x-z<x. Applying (4.1) and (4.2) of Lemma 4.1, we have
 $\sum_{\begin{subarray}{c}1\leq k\leq x\\ \sqrt{y}<P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\left(\sum_{\begin{subarray}{c}1\leq k\leq x\\ P^{+}(k)\leq y\end{subarray}}\frac{\Lambda(k)}{k}\omega_{f}(k)\right)^{n-1}$
@@
 Let $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 2$.
-Put $h=x-z$.
+Put h=xzh=x-z.
 Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
@@
 $W_{m}^{+}$ $\leq\frac{h}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}$
-×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
+$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~(j=1,2)\\ P^{+}(k_{j})\leq y~(j\geq 3)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
 $\quad+\frac{1}{(\log{f(z)})^{2}\left(\log{\frac{f(z)}{x}}\right)^{m-1}}$
-×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
+$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
 by the definition of $W_{m}^{+}$. We have
@@
 Then we divide the sums into those over $k_{1},\ldots,k_{m+1}\geq 1$ such that $(k_{j},\Delta_{f})=1$ if $j\in S$ and $(k_{j},\Delta_{f})>1$ otherwise, where $S$ is a subset of $\{1,\ldots,m+1\}$.
-Recall that $[k_{1},k_{2}]k_{3}\cdots k_{m+1}=k_{1}k_{2}k_{3}\cdots k_{m+1}$ holds if $(k_{1},k_{2})=1$.
+Recall that $k_{1},k_{2}]k_{3}\cdots k_{m+1}=k_{1}k_{2}k_{3}\cdots k_{m+1}$ holds if $(k_{1},k_{2})=1$.
 Therefore we obtain
@@
 $<\left(\frac{1}{2}+\frac{1}{2}\frac{2g+1}{8du}+\sqrt{\frac{1}{2}\frac{2g+1}{8du}+\frac{1}{4}\left(\frac{2g+1}{8du}\right)^{2}}\right)$
-×$\qquad\times\frac{xg^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right)$
+$\qquad\times\frac{xg^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right)$
 $=\gamma_{f}(u)\frac{g^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(x+O\left(\frac{x}{\log\log{x}}\right)\right).$
 We have trivially $\Psi_{f}(z,x^{1/u})\leq z$. Hence the desired bound for $\Psi_{f}(x,x^{1/u})$ follows by the choice of $z$ as in (4.24). ∎
-5. Proof of Theorem 1.2
-Let $\alpha$ be an arbitrary algebraic irrational number, and $K=\mathbb{Q}(\alpha)$ the algebraic field generated by $\alpha$ over the rationals.
-Then we define $P_{K}$ as the set of all prime ideals $\mathfrak{p}$ of $K$ such that the following conditions are satisfied:
+5. Proof of Theorem 1.2 and its corollary
+Let $\alpha$ be an arbitrary algebraic irrational number, and $K=\mathbb{Q}(\alpha)$ the algebraic field generated by $\alpha$ over the rationals. For any rational integer $n\geq 0$, we have
+$(n+\alpha)\mathfrak{a}=\prod_{\mathfrak{p}\in J_{\alpha}}\mathfrak{p}^{u_{n}(\mathfrak{p})},$
+where $J_{\alpha}$ denotes the set of all prime ideals of $K$, and $u_{n}(\mathfrak{p})$ are non-negative integers.
+Define $P_{\alpha}$ as the subset of $J_{\alpha}$ consisting of prime ideals $\mathfrak{p}$ such that
   • $(\mathrm{i})$
@@
     $\mathfrak{p}$ is unambiguous.
-In other words, we have $\operatorname{N}(\mathfrak{p})=p$ and $\mathfrak{p}^{2}\nmid(p)$ for a rational prime $p$ if $\mathfrak{p}\in P_{K}$.
-Let $\mathfrak{a}$ be the ideal denominator of the fractional ideal $(\alpha)$ of $K$.
-Then we factorize the ideal $(n+\alpha)\mathfrak{a}$ for $n\in\mathbb{Z}$ as $(n+\alpha)\mathfrak{a}=\mathfrak{a}_{n}\mathfrak{b}_{n}$, where
-$\mathfrak{a}_{n}=\prod_{\mathfrak{p}\in P_{K}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}\quad\text{and}\quad\mathfrak{b}_{n}=\prod_{\mathfrak{p}\notin P_{K}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)},$
-with $v_{\mathfrak{p}}(n+\alpha)\in\mathbb{Z}_{\geq 0}$ such that $v_{\mathfrak{p}}(n+\alpha)=0$ for all but finitely many of the prime ideals $\mathfrak{p}$.
-We begin by proving the following preliminary lemmas.
-Lemma 5.1.
-There exists a constant $C(\alpha)>0$ depending only on $\alpha$ such that
-$\operatorname{N}(\mathfrak{b}_{n})\leq C(\alpha)$
+In other words, we have $\operatorname{N}(\mathfrak{p})=p$ and $\mathfrak{p}^{2}\nmid(p)$ for a rational prime $p$ if $\mathfrak{p}\in P_{\alpha}$.
+Then we rewrite the factorization of $(n+\alpha)\mathfrak{a}$ as
+$(n+\alpha)\mathfrak{a}=\mathfrak{b}_{n}\prod_{\mathfrak{p}\in P_{\alpha}}\mathfrak{p}^{u_{n}(\mathfrak{p})},$
+where $\mathfrak{b}_{n}$ is the integral containing all prime factors of $(n+\alpha)\mathfrak{a}$ which are not in $P_{\alpha}$.
+By [16, Lemma 4.4], we know that there exists a constant $C(\alpha)>0$ depending at most on $\alpha$ such that
+(5.1) $\operatorname{N}(\mathfrak{b}_{n})\leq C(\alpha)$
 holds for any $n\in\mathbb{Z}$.
-Proof.
-Take a prime ideal $\mathfrak{p}$ of $K$ and $v\geq 1$ such that $\mathfrak{p}^{v}\mid\mathfrak{b}_{n}$.
-We consider an upper bound for the norm of $\mathfrak{p}^{v}$.
-Let $L$ denote the Galois closure of $K$ over $\mathbb{Q}$.
-For a rational prime $p$ with $\mathfrak{p}\mid p$, we write the factorization of $p$ in $L$ as
-$p\mathcal{O}_{L}=\left(\mathfrak{P}_{1}\cdots\mathfrak{P}_{g}\right)^{e},$
-where $\mathfrak{P}_{1},\ldots,\mathfrak{P}_{g}$ are prime ideals of $L$ with $N(\mathfrak{P}_{k})=p^{f}$.
-Note that at least one of $e\geq 2$ and $f\geq 2$ is true.
-Indeed, if $e=f=1$, then $p$ splits completely in $L$.
-This implies that $p$ splits completely also in $K$, but $\mathfrak{p}\in P_{K}$ follows in this case.
-Next, we factorize the prime ideal $\mathfrak{p}$ as
-(5.1) $\mathfrak{p}\mathcal{O}_{L}=\left(\mathfrak{P}_{k_{1}}\cdots\mathfrak{P}_{k_{m}}\right)^{\eta},$
-where $k_{1},\ldots,k_{m}\in\{1,\ldots,g\}$ and $1\leq\eta\leq e$.
-Denote by $M_{j}$ the decomposition field of $\mathfrak{P}_{k_{j}}$ over $\mathbb{Q}$, and define $\mathfrak{q}_{j}=\mathfrak{P}_{k_{j}}\cap\mathcal{O}_{M_{j}}$.
-Then $\mathfrak{P}_{k_{j}}$ is the unique prime ideal of $L$ lying over $\mathfrak{q}_{j}$.
-More precisely, we see that
-(5.2) $\mathfrak{q}_{j}\mathcal{O}_{L}=\mathfrak{P}_{k_{j}}^{e}\quad\text{and}\quad\operatorname{N}(\mathfrak{q}_{j})=p$
-by the standard properties of decomposition fields. Put $v\eta=eq+r$ with $q\geq 0$ and $0\leq r<e$. From (5.1) and (5.2), we obtain
-(5.3) $\mathfrak{p}^{v}\mathcal{O}_{L}$ $=\left(\mathfrak{P}_{k_{1}}\cdots\mathfrak{P}_{k_{m}}\right)^{v\eta}$
-$=\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}\cdot\mathfrak{P}_{k_{1}}^{r}\cdots\mathfrak{P}_{k_{m}}^{r}.$
-Hence, taking the norms, we have
-(5.4) $\operatorname{N}(\mathfrak{p}^{v})^{[L:K]}=\operatorname{N}(\mathfrak{q}_{1}^{q})^{d}\cdots\operatorname{N}(\mathfrak{q}_{m}^{q})^{d}\cdot\operatorname{N}(\mathfrak{P}_{k_{1}})^{r}\cdots\operatorname{N}(\mathfrak{P}_{k_{m}})^{r},$
-where $d=[L:M_{j}]=ef$.
-Denote by $\Delta_{L/\mathbb{Q}}$ the absolute value of the discriminant of $L$ over $\mathbb{Q}$.
-If $e\geq 2$, then $p\mid\Delta_{L/\mathbb{Q}}$ is satisfied.
-In this case, $\operatorname{N}(\mathfrak{P}_{k_{j}})^{r}=p^{fr}\leq\Delta_{L/\mathbb{Q}}^{ef}$ holds.
-If $e=1$, then $\operatorname{N}(\mathfrak{P}_{k_{j}})^{r}=1$ since $r$ must be equal to 00.
-As a result, we obtain
-(5.5) $\operatorname{N}(\mathfrak{P}_{k_{1}})^{r}\cdots\operatorname{N}(\mathfrak{P}_{k_{m}})^{r}\leq\Delta_{L/\mathbb{Q}}^{efm}\leq\Delta_{L/\mathbb{Q}}^{[L:\mathbb{Q}]}$
-due to $[L:\mathbb{Q}]=efg$.
-Next, we choose $\beta\in\mathcal{O}_{K}$ and $c\in\mathbb{Z}_{\geq 1}$ so that $\alpha=\beta/c$.
-Note that we can determine them uniquely from $\alpha$ by fixing $c$ as the minimum one with this property.
-Recall that $\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}$ divides the ideal $(n+\alpha)\mathfrak{a}\mathcal{O}_{L}$ by (5.3) due to $\mathfrak{p}^{v}\mid\mathfrak{b}_{n}$.
-Then we obtain
-$\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}\mid(nc+\beta)\mathcal{O}_{L}$
-since $(n+\alpha)\mathfrak{a}$ divides $(nc+\beta)$.
-In other words, $nc+\beta$ belongs to $\mathfrak{q}_{j}^{q}\mathcal{O}_{L}$ for any $j$.
-Suppose that $\mathcal{O}_{L}=\mathcal{O}_{M_{j}}[x_{1},\ldots,x_{d}]$ with $x_{1},\ldots,x_{d}\in\mathcal{O}_{L}$.
-Then we have
-$1=a_{1}x_{1}+\cdots+a_{d}x_{d}\quad\text{and}\quad\beta=b_{1}x_{1}+\cdots+b_{d}x_{d}$
-with some constants $a_{\ell},b_{\ell}\in\mathcal{O}_{M_{j}}$. It yields the formula
-$nc+\beta=(na_{1}c+b_{1})x_{1}+\cdots+(na_{d}c+b_{d})x_{d}.$
-Then $nc+\beta\in\mathfrak{q}_{j}^{q}\mathcal{O}_{L}$ is equivalent that
-$(na_{1}c+b_{1})x_{1}+\cdots+(na_{d}c+b_{d})x_{d}=\sum_{t=1}^{T}p_{t}z_{t}$
-with some $p_{t}\in\mathfrak{q}_{j}^{q}$ and $z_{t}\in\mathcal{O}_{L}$. We have
-$\sum_{t=1}^{T}p_{t}z_{t}=\sum_{t=1}^{T}p_{t}(A_{t1}x_{1}+\cdots+A_{td}x_{d}),$
-where $A_{t1},\ldots,A_{td}\in\mathcal{O}_{M_{j}}$. Comparing them, we derive
-$na_{\ell}c+b_{\ell}=\sum_{t=1}^{T}p_{t}A_{t\ell}\in\mathfrak{q}_{j}^{q}$
-for any $\ell\in\{1,\ldots,d\}$.
-This implies that $a_{k}b_{\ell}-a_{\ell}b_{k}\in\mathfrak{q}_{j}^{q}$ for any $k,\ell\in\{1,\ldots,d\}$.
-Hence we conclude that $\mathfrak{q}_{j}^{q}$ divides the integral ideal generated by all elements in $\{a_{k}b_{\ell}-a_{\ell}b_{k}\mid k\neq\ell\}$ over $\mathcal{O}_{M_{j}}$.
-Define $G(\alpha;M_{j})$ as the norm of this ideal.
-One can check that $G(\alpha;M_{j})$ is independent to the choice of the basis $\{x_{1},\ldots,x_{d}\}$.
-Furthermore, $G(\alpha)=\max G(\alpha;M)$ depends only on $\alpha$, where the maximum is taken over all algebraic fields $M$ such that $\mathbb{Q}\subset M\subsetneq L$.
-Finally, we see that
-(5.6) $\operatorname{N}(\mathfrak{q}_{1}^{q})^{d}\cdots\operatorname{N}(\mathfrak{q}_{m}^{q})^{d}\leq G(\alpha)^{dm}\leq G(\alpha)^{[L:\mathbb{Q}]}$
-due to $d=ef$ and $[L:\mathbb{Q}]=efg$. By (5.4), (5.5), and (5.6), we arrive at
-$\operatorname{N}(\mathfrak{p}^{v})\leq\left(G(\alpha)\Delta_{L/\mathbb{Q}}\right)^{[K:\mathbb{Q}]}=C_{1}(\alpha),$
-say. Then the desired result immediately follows. Indeed, we have
-$\operatorname{N}(\mathfrak{b}_{n})=\prod_{\mathfrak{p}\mid\mathfrak{b}_{n}}\operatorname{N}(\mathfrak{p}^{\nu_{\mathfrak{p}}(\mathfrak{b}_{n})})\leq C_{1}(\alpha)^{\pi_{K}(C_{1}(\alpha))},$
-where $\pi_{K}(x)$ denotes the number of prime ideals $\mathfrak{p}$ of $K$ such that $\operatorname{N}(\mathfrak{p})\leq x$. ∎
+Thus we have $\operatorname{N}(\mathfrak{p})\leq C(\alpha)$ for any $\mathfrak{p}\in J_{\alpha}\setminus P_{\alpha}$.
+Then we prove Theorem 1.2 by using the following lemmas.
 Lemma 5.1.
 Take an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$, and denote by $A_{f}$ its leading coefficient.
-Let $C(\alpha)>0$ be the constant described in Lemma 5.1.
+Let $C(\alpha)>0$ be the constant described in (5.1).
 Then, for any $n\in\mathbb{Z}$ and any $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$, the following conditions are equivalent.
@@
   • $(\mathrm{B})$
-    There exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
+    There exists a prime ideal $\mathfrak{p}\in P_{\alpha}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
 Proof.
@@
 Thus we have that $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$.
-Furthermore, it implies that there exists a prime ideal $\mathfrak{p}$ of $K$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})=p^{f}$ with some $j\geq 1$.
-If we suppose that $\mathfrak{p}\notin P_{K}$, then $\operatorname{N}(\mathfrak{p})>x>C(\alpha)$ would follow, which yields contradiction by Lemma 5.1.
-Hence we conclude that $\mathfrak{p}\in P_{K}$ holds, and therefore, $(\mathrm{A})$ implies $(\mathrm{B})$.
-Suppose conversely that there exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
-Then $\operatorname{N}(\mathfrak{p})=p$ holds for a rational prime $p$ by the definition of $P_{K}$.
+Furthermore, it implies that there exists a prime ideal $\mathfrak{p}$ of $K$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})=p^{f}$ with some $f\geq 1$.
+If we suppose that $\mathfrak{p}\notin P_{\alpha}$, then $\operatorname{N}(\mathfrak{p})>x>C(\alpha)$ would follow, which yields contradiction.
+Hence we conclude that $\mathfrak{p}\in P_{\alpha}$ holds, and therefore, $(\mathrm{A})$ implies $(\mathrm{B})$.
+Suppose conversely that there exists a prime ideal $\mathfrak{p}\in P_{\alpha}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
+Then $\operatorname{N}(\mathfrak{p})=p$ holds for a rational prime $p$ by the definition of $P_{\alpha}$.
 Thus $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$ and $p>x$ are satisfied.
@@
 ∎
-Lemma 5.3.
-Let $\mathfrak{p}\in P_{K}$ and $v\geq 1$.
+Lemma 5.2 (Lemma 4.6 in [16]).
+Let $\mathfrak{p}\in P_{\alpha}$ and $v\geq 1$.
 Assume $\mathfrak{p}^{v}\mid(m+\alpha)\mathfrak{a}$ and $\mathfrak{p}^{v}\mid(n+\alpha)\mathfrak{a}$ for $m,n\in\mathbb{Z}$.
 Then we have $m\equiv n~(\bmod\,{p^{v}})$ with $\operatorname{N}(\mathfrak{p})=p$.
-Proof.
-Put $m-n=p^{w}A$ with some $w\in\mathbb{Z}_{\geq 0}$ and $A\in\mathbb{Z}_{\geq 1}$ such that $(p,A)=1$.
-By the assumption, we have $(m+\alpha)\mathfrak{a}\subset\mathfrak{p}^{v}$ and $(n+\alpha)\mathfrak{a}\subset\mathfrak{p}^{v}$.
-Thus we see that $(m-n)\mathfrak{a}\subset\mathfrak{p}^{v}$ holds, which yields
-$\mathfrak{p}^{v}\mid(m-n)$
-due to $\mathfrak{p}\nmid\mathfrak{a}$. Therefore, $\mathfrak{p}^{v}\mid(p)^{w}(A)$ follows. We deduce that $\mathfrak{p}\nmid(A)$ from $(p,A)=1$ by taking the norms. As a result, we obtain
-$\mathfrak{p}^{v}\mid(p)^{w}.$
-Recall that $\mathfrak{p}$ is unambiguous, i.e. $\mathfrak{p}^{2}\nmid(p)$. Hence we see that $w\geq v$, which implies $p^{v}\mid(m-n)$ as desired. ∎
-From the above, we associate $C_{\alpha}(x)$ defined in Section 1.2 with $\Psi_{f}(x,y)$ for an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$.
-The relation between $C_{\alpha}(x)$ and $\Psi_{f}(x,y)$ was mentioned firstly by Worley [20], and a more detailed research was presented by Andersson in his unpublished work [1].
-In this paper, we obtain the following result.
-Proposition 5.4.
-Take an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$. Then the asymptotic formula
-$C_{\alpha}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x}{\log{x}}\right)$
-holds for sufficiently large $x$, where the implied constant depends only on $K=\mathbb{Q}(\alpha)$.
-Proof.
+Proof of Theorem 1.2.
 Denote by $M_{\alpha}(x)$ the set of integers $n$ in $[1,x]$ such that there exists a prime ideal $\mathfrak{p}$ of $K$ with conditions (1.5).
@@
 Here, we suppose $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$ as in Lemma 5.1.
-Then we deduce from Lemma 5.2 that there exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
+Then we deduce from Lemma 5.1 that there exists a prime ideal $\mathfrak{p}\in P_{\alpha}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
 If $\mathfrak{p}\mid(m+\alpha)\mathfrak{a}$ is satisfied for some $m\in\mathbb{Z}$ in $[1,x]$, then we apply Lemma 5.2 to deduce that $m\equiv n~(\bmod\,{p})$ with $\operatorname{N}(\mathfrak{p})=p$.
@@
 by the prime ideal theorem, where the implied constant depends only on $K$. It yields the desired result by formulas (5.3). ∎
-Proof of Theorem 1.2.
-Let $f(t)\in\mathbb{Z}[t]$ be an irreducible polynomial with $f(-\alpha)=0$. Then $f(t)$ is of degree $d=\deg(\alpha)\geq 2$. By Proposition 5.4 and Theorem 1.1 with $u=1$, we have
+Proof of Corollary 1.3.
+Let $f(t)\in\mathbb{Z}[t]$ be an irreducible polynomial with $f(-\alpha)=0$. Then $f(t)$ is of degree $d=\deg(\alpha)\geq 2$. By Theorem 1.1 with $u=1$, we deduce from Theorem 1.2 that
 $C_{\alpha}(x)>x-\frac{\gamma_{f}(1)}{d}\left(x+O\left(\frac{x}{\log\log{x}}\right)\right)+O\left(\frac{x}{\log{x}}\right),$
@@
 by using (1.3). We also obtain
-$0.543\displaystyle 1-\frac{\gamma_{f}(1)}{d}>1-\frac{0.914}{2}=$ 0.543
+$1-\frac{\gamma_{f}(1)}{d}>1-\frac{0.914}{2}=0.543$
 by (1.4) and $d\geq 2$. Hence the proof is completed. ∎
-Let $f(t)$ be an irreducible polynomial with $f(-\alpha)=0$. If (1.1) is true at least for $u=1$, then we deduce from Proposition 5.4 that
+Let $f(t)$ be an irreducible polynomial with $f(-\alpha)=1$. If (1.1) is true at least for $u=1$, then we deduce from Theorem 1.2 that
 $C_{\alpha}(x)\sim(1-\rho(d))x$
 as $x\to\infty$, where we put $d=\deg(\alpha)\geq 2$. Here, it is known that $\rho(d)=d^{-d+o(d)}$ as $d\to\infty$. Thus the result of Theorem 1.2 still leaves room for improvement.
-Remark 5.5.
-Actually, Lemma 5.1 is much stronger than necessary for the proof of Theorem 1.2 since one needs just that $\operatorname{N}(\mathfrak{p})$ is bounded for any $\mathfrak{p}\notin P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$.
-We note that Lemma 5.1 is a modification of a statement of Cassels [2].
-He used the set $P_{K}^{*}$ consisting of prime ideals $\mathfrak{p}$ of $K$ such that
-  • $(\mathrm{i})$
-    $\mathfrak{p}$ is of the first degree;
-  • $(\mathrm{ii})$
-    $\mathfrak{p}$ is unambiguous;
-  • $(\mathrm{iii})$
-    for any $m\in\mathbb{Z}$, if $\mathfrak{p}\mid(m+\alpha)\mathfrak{a}$ then $(m+\alpha)\mathfrak{a}$ is not divisible by any prime ideal $\mathfrak{p}^{\prime}\neq\mathfrak{p}$ satisfying $\operatorname{N}(\mathfrak{p})=\operatorname{N}(\mathfrak{p}^{\prime})$.
-We factorize the ideal $(n+\alpha)\mathfrak{a}$ as $(n+\alpha)\mathfrak{a}=\mathfrak{a}_{n}^{*}\mathfrak{b}_{n}^{*}$, where
-$\mathfrak{a}_{n}^{*}=\prod_{\mathfrak{p}\in P_{K}^{*}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}\quad\text{and}\quad\mathfrak{b}_{n}^{*}=\prod_{\mathfrak{p}\notin P_{K}^{*}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}.$
-Then it was claimed in [2, p.
-181] that $\operatorname{N}(\mathfrak{b}_{n}^{*})$ is bounded for $n\in\mathbb{Z}$.
-However, we know that $\operatorname{N}(\mathfrak{b}_{n}^{*})$ is not bounded in general.
-For example, letting $\alpha=7\sqrt{2}$, one can find for any $v\geq 1$ an integer $n=n(v)$ such that $\operatorname{N}(b_{n}^{*})\geq 7^{v}$.
-See [16, Remark 1.6] for more information.
-Hence the original method by Cassels seems flawed.
-The boundedness of $\operatorname{N}(\mathfrak{b}_{n}^{*})$ was used in [2] to deduce the inequality
-$\sum_{\mathfrak{p}\in P_{K}^{*}}v_{\mathfrak{p}}(n+\alpha)\log{\operatorname{N}(\mathfrak{p})}>2\log{n}-C$
-for a constant $C>0$ which is independent of $n$.
-Furthermore, condition $(\mathrm{iii})$ was used to see that there is no contribution of terms for $\mathfrak{p}_{1},\mathfrak{p}_{2}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})$ in the sum
-(5.9) $\sum_{n\in\mathfrak{S}_{N}}\sum_{\begin{subarray}{c}\mathfrak{p}_{1},\mathfrak{p}_{2}\in P_{K}^{*}\\ \mathfrak{p}_{j}\mid(n+\alpha)\mathfrak{a}\\ \operatorname{N}(\mathfrak{p}_{j})<\sqrt{M}\end{subarray}}\frac{\log{\operatorname{N}(\mathfrak{p}_{1})}}{\operatorname{N}(\mathfrak{p}_{1})}\frac{\log{\operatorname{N}(\mathfrak{p}_{2})}}{\operatorname{N}(\mathfrak{p}_{2})},$
-where $\mathfrak{S}_{N}$ is the set of integers $n$ in $[N,N+M]$ such that $\operatorname{N}(\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)})<M$ holds for all $\mathfrak{p}\in P_{K}^{*}$ with $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$.
-Note that (5.9) appeared during a calculation using the Cauchy–Schwarz inequality, and that a similar argument was also presented in the proof of Proposition 2.1.
-The difference is that the method of this paper is non ideal-theoretic, which requires nothing similar to condition $(\mathrm{iii})$.
-On the other hand, it is also possible to recover the proof in [2] by replacing $P_{K}^{*}$ with $P_{K}$.
-Indeed, $\operatorname{N}(\mathfrak{b}_{n})$ is bounded by Lemma 5.1, which yields
-$\sum_{\mathfrak{p}\in P_{K}}v_{\mathfrak{p}}(n+\alpha)\log{\operatorname{N}(\mathfrak{p})}>2\log{n}-C.$
-Furthermore, for any prime ideals $\mathfrak{p}_{1},\mathfrak{p}_{2}\in P_{K}$ with $\mathfrak{p}_{j}\mid(n+\alpha)\mathfrak{a}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})=:p$, one can show that $p$ is bounded for $n\in\mathbb{Z}$.
-Actually, this is sufficient to show that the contribution of terms for $\mathfrak{p}_{1},\mathfrak{p}_{2}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})$ in (5.9) is not zero but can be negligible.
-To close this section, we confirm the original statement of Cassels [2].
+Remark that the formulation of Corollary 1.3 is due to Worley [21], and the original statement by Cassels in [2] is slightly different.
+To close this section, we confirm the original statement of Cassels.
 Theorem 5.3 (Cassels [2]).
@@
 Suppose that $N\geq N_{0}$ and put $M=\lfloor 10^{-6}N\rfloor$.
-Then at least $10051M/$ 100 integers in $N<n\leq N+M$ are such that $(n+\alpha)\mathfrak{a}$ is divisible by a prime ideal $\mathfrak{p}$ which does not divide $(m+\alpha)\mathfrak{a}$ for any integer $0\leq m\leq N+M$ with $m\neq n$.
+Then at least $51M/100$ integers in $N<n\leq N+M$ are such that $(n+\alpha)\mathfrak{a}$ is divisible by a prime ideal $\mathfrak{p}$ which does not divide $(m+\alpha)\mathfrak{a}$ for any integer $0\leq m\leq N+M$ with $m\neq n$.
 Proof.
@@
 $\Psi_{f}(N+M,N+M)-\Psi_{f}(N,N+M)<\frac{\kappa_{f}(1)}{d}M<0.48M$
-by (2.2) if $N$ is large enough. Furthermore, along the same line as the proof of Proposition 5.4, we see that the number of integers in $N<n\leq N+M$ with the desired property is equal to
+by (2.2) if $N$ is large enough. Furthermore, along the same line as the proof of Theorem 1.2, we see that the number of integers in $N<n\leq N+M$ with the desired property is equal to
 $M-\left(\Psi_{f}(N+M,N+M)-\Psi_{f}(N,N+M)\right)+O\left(\frac{N+M}{\log(N+M)}\right),$
 where $f(t)\in\mathbb{Z}[t]$ is an irreducible polynomial with $f(-\alpha)=0$. Hence we obtain the conclusion. ∎
-6. Proof of Theorems 1.3 and 1.4
-We prove Theorems 1.3 and 1.4 by associating the functions $R_{b}(x)$ and $N(x)$ defined in Section 1.3 with $\Psi_{f}(x,y)$ for a suitable polynomial $f(t)$.
-For this, we recall the following criterions.
+6. Proof of Theorem 1.4 and its corollaries
+Firstly, we recall the following criterions, which yield $N(x)=R_{1}(x)$ immediately.
 Lemma 6.1 (Everest–Stevens–Tamsett–Ward [7]).
-Let $n\in\mathbb{Z}_{>|b|}$ with $-b$ not an integer square. Then $n^{2}+b$ has a primitive divisor if and only if $P^{+}(n^{2}+b)>2n$.
+Let $n\in\mathbb{Z}_{>|b|}$ with b-b not an integer square. Then $n^{2}+b$ has a primitive divisor if and only if $P^{+}(n^{2}+b)>2n$.
 Lemma 6.2 (Todd [20]).
 Let $n\in\mathbb{Z}_{\geq 1}$. Then $\arctan n$ is irreducible if and only if $P^{+}(n^{2}+1)>2n$.
-Then the equality $N(x)=R_{1}(x)$ immediately follows. Furthermore, we associate $R_{b}(x)$ with $\Psi_{f}(x,y)$ for the quadratic polynomial $f(t)=t^{2}+b$ as follows.
-Proposition 6.3.
-Let $f(t)=t^{2}+b$ with $-b$ not an integer square. Then we have
-$R_{b}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x\log\log{x}}{\log{x}}\right)$
-for sufficiently large $x$, where the implied constant depends only on $|b|$.
-Proof.
+Furthermore, Theorem 1.4 is proved by using Lemma 6.1 as follows.
+Proof of Theorem 1.4.
 By Lemma 6.1, $R_{b}(x)$ is evaluated as
@@
 $S_{4}$ $\leq\sum_{\begin{subarray}{c}z<n\leq x\\ 2z<P^{+}(n^{2}+b)<2x\end{subarray}}1$
-with a parameter $z$ such that $21\leq z\leq x/$ 2. From the above, we obtain the formula
+with a parameter $z$ such that $1\leq z\leq x/2$. From the above, we obtain the formula
 (6.1) $R_{b}(x)=x-\Psi_{f}(x,x)+O\left(z+\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1\right)$
 for $x\geq 1$. Here, we see that
-$\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1=\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ P^{+}(n^{2}+b)=p\end{subarray}}1\leq\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0~{}\pmod{p}\end{subarray}}1$
+$\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1=\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ P^{+}(n^{2}+b)=p\end{subarray}}1\leq\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0~\hskip-5.69054pt\pmod{p}\end{subarray}}1$
 holds since $P^{+}(n^{2}+b)=p$ implies $f(n)\equiv 0~(\bmod\,{p})$. We have $\omega_{f}(p)\leq 4\sqrt{|b|}$ by applying (4.5). Therefore, the last sum is further evaluated as
-$\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0\pmod{p}\end{subarray}}1$ $\ll\sum_{2z\leq p<2x}\left(\frac{x}{p}+1\right)$
+$\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0\hskip-5.69054pt\pmod{p}\end{subarray}}1$ $\ll\sum_{2z\leq p<2x}\left(\frac{x}{p}+1\right)$
 $=x\log\left(\frac{\log(2x)}{\log(2z)}\right)+O\left(\frac{x}{\log(2z)}\right)$
@@
 which yields the desired result by (6.1). ∎
-If (1.1) is true for $f(t)=t^{2}+b$ and $u=1$, then Proposition 6.3 yields that
-$R_{b}(x)\sim(1-\rho(2))x$
-as $x\to\infty$. Recall that $\rho(u)=1-\log{u}$ for $1\leq u\leq 2$. Hence (1.1) contains the conjectures of Everest–Harman and Chowla–Todd on $R_{b}(x)$ and $N(x)$, respectively.
-Proof of Theorems 1.3 and 1.4.
-The proof is similar to that of Theorem 1.2. Let $f(t)=t^{2}+b$. By Proposition 6.3 and Theorem 1.1 with $u=1$, we have
+Proof of Corollaries 1.5 and 1.6.
+The proof is similar to that of Corollary 1.3. Let $f(t)=t^{2}+b$. By Theorem 1.1 with $u=1$, we deduce from Theorem 1.4 that
 $R_{b}(x)>x-\frac{\gamma_{f}(1)}{2}\left(x+O\left(\frac{x}{\log\log{x}}\right)\right)+O\left(\frac{x}{\log{x}}\right),$
@@
 for sufficiently large $x$. Since we have
-$0.543\displaystyle 1-\frac{\gamma_{f}(1)}{2}>1-\frac{0.914}{2}=$ 0.543
+$1-\frac{\gamma_{f}(1)}{2}>1-\frac{0.914}{2}=0.543$
 by using (1.4), we obtain the result for $R_{b}(x)$. The result for $N(x)$ also follows due to $N(x)=R_{1}(x)$ as explained before. ∎